### v2

# New versions of Frobenius and integral closure of ideals

Abstract.
In this article, we define two new versions of integral and Frobenius closures of ideals which incorporate an auxiliary ideal and a real parameter. These additional ingredients are common in adjusting old definitions of ideal closures in order to generalize them to pairs, with an eye towards further applications in algebraic geometry. In the case of tight closure, similar generalizations exist due to N. Hara and K.I. Yoshida, as well as A. Vraciu. We study their basic properties and give computationally effective calculations of the adjusted tight, Frobenius, and integral closures in the case of affine semigroup rings in terms of the convex geometry of the associated exponent sets.
2020 Mathematics Subject Classification. Primary 13A35, 13B22, Secondary 13D40
Keyword: Integral closure, tight closure, Frobenius closure, semigroup rings, joint Hilbert-Kunz multiplicity
1. Introduction
The tight closure of an ideal was introduced by M. Hochster and C. Huneke in [HH90], and has since been indispensable in the study of rings of prime characteristic. Tight closure even illuminates properties of rings of characteristic zero via reduction to prime characteristic techniques. In particular, deep uniformity theorems such as the Briançon-Skoda theorem (see the discussion around [HH90, Theorem 5.4]) eluded proofs with algebraic techniques in characteristic zero, but have relatively simple proofs in the language of tight closure.
In studying tight closure, one is naturally led to study test ideals, which uniformly witness tight closure relations for every ideal in the ring simultaneously. N. Hara [Har01] and K. E. Smith [Smi00] showed that test ideals in prime characteristic correspond in a natural way to multiplier ideals in birational geometry, and so are deeply related to geometric concepts like resolutions of singularities. The initial connection made by Hara and Smith was to the multiplier ideal of the trivial pair, but a strength of the multiplier ideal concept is that one may vary a $\mathbb{Q}$-divisor or an ideal of the ring and obtain information about the singularities in the locus of primes containing that ideal. For a survey of test ideals and their connections to multiplier ideals, we recommend [ST12].
Hara and K. Yoshida [HY03] introduced an adjusted form of tight closure, called $\mathfrak{a}^{t}$-tight closure (see 3.9 below), which incorporates an auxiliary ideal and a non-negative real parameter. The benefit to this adjustment is that now the appropriate notion of test ideals for $\mathfrak{a}^{t}$-tight closure corresponds to the multiplier ideal of the ideal pair $(R,\mathfrak{a}^{t})$. One disadvantage of the Hara-Yoshida tight closure is that it is not a true closure operation, in that the operation is not idempotent.
To rectify this issue, A. Vraciu [Vra08] introduced a new version of $\mathfrak{a}^{t}$-tight closure, which is a true ideal closure operation. Her version is contained in the Hara-Yoshida $\mathfrak{a}^{t}$-tight closure, and captures subtly different information. In some cases the test ideals of the Hara-Yoshida and Vraciu versions agree, see [Vra08, Theorem 4.3]. Similar versions of adjusted tight and (sharp) Frobenius closures are studied in [STV17], where the authors use these closures to define and characterize $F$-rationality and sharp $F$-injectivity of the pair $(R,\mathfrak{a}^{t}),$ and also to define the ($F$-pure) threshold of the ideal $\mathfrak{a}.$
Our aim in this article is to repeat the adjustments of tight closure made by Hara-Yoshida and Vraciu to Frobenius and integral closure, which are two more classical closures near to the tight closure.
Definition (3.1, 3.12).
Fix a real number $t>0$, and let $I$ and $\mathfrak{a}$ be ideals in a ring $R$. Let $R^{\circ}$ all elements of $R$ not in a minimal prime of $R$. For the case of Frobenius closure, we assume that $R$ is of prime characteristic $p>0$.
  1. (a)
    The Hara-Yoshida $\mathfrak{a}^{t}$-Integral closure of $I$, denoted $\overline{I}^{\mathfrak{a}^{t}}$, is defined by
    $\overline{I}^{\mathfrak{a}^{t}}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{ and all }n\gg 0,\,cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}\right.\right\}.$
  2. (b)
    The Vraciu $\mathfrak{a}^{t}$-Integral closure of $I$, denoted ${}^{\mathfrak{a}^{t}}\overline{I}$, is defined by
    ${}^{\mathfrak{a}^{t}}\overline{I}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{ and all }n\gg 0,\,cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset\mathfrak{a}^{\lceil tn\rceil}I^{n}\right.\right\}.$
  3. (c)
    The Hara-Yoshida $\mathfrak{a}^{t}$-Frobenius closure of $I$, denoted $I^{F\mathfrak{a}^{t}},$ is defined by
    $I^{F\mathfrak{a}^{t}}=\left\{x\in R\left|\text{ for all }e\gg 0,\,x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{\left[p^{e}\right]}\right.\right\}.$
  4. (d)
    The Vraciu $\mathfrak{a}^{t}$-Frobenius closure of $I$, denoted ${}^{\mathfrak{a}^{t}}I^{F}$, is defined by
    ${}^{\mathfrak{a}^{t}}I^{F}=\left\{x\in R\left|\text{ for all }e\gg 0,\,x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset\mathfrak{a}^{\lceil tp^{e}\rceil}I^{\left[p^{e}\right]}\right.\right\}.$
Throughout Section 3, we show that the properties of these new adjusted closures are similar to the adjusted tight closures and their usual closures. In Subsection 3.4, we compare the new closures to each other and give an adjusted version of the Briançon-Skoda theorem.
Theorem (Theorem 3.19).
Let $\mathfrak{a},I$ be ideals in a ring $R$ of prime characteristic $p>0$ where $I\cap R^{\circ}$ and $\mathfrak{a}\cap R^{\circ}$ are non-empty. Let $I$ be generated by $d$ elements and let $t\in\mathbb{R}_{\geq 0}$.
  1. (a)
    Then $\overline{I^{d+m}}^{\mathfrak{a}^{t}}\subset$ $(I^{m+1})^{*\mathfrak{a}^{t}}$ for all $m\geq 0,$ and
  2. (b)
    $\overline{I^{d+m}}^{\mathfrak{a}^{t}}\subset(I^{m})^{F\mathfrak{a}^{t}},$ for all $m\geq 1.$
One unfortunate aspect of the adjusted closure operations are that they are even more difficult to compute than their usual counterparts. However, in some cases, they are related to colon ideals of the usual closure operation. Intriguingly, we also relate the adjusted tight and integral closure to the rational powers $\mathfrak{a}_{t}$ of an ideal (see 2.4).
Theorem (3.5, 3.10, Proposition 3.16).
Let $R$ be a ring and $I,\mathfrak{a}\subset R$ be ideals such that $I\cap R^{\circ}$ and $\mathfrak{a}\cap R^{\circ}$ are non-empty. Further, let $t\geq 1$ be a rational number. Then,
  1. (a)
    $\overline{I}^{\mathfrak{a}^{t}}\subset\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)$ with equality if $\mathfrak{a}$ is principal or if $t$ is an integer, and
  2. (b)
    $\overline{I}^{\mathfrak{a}^{t}}\subset(\overline{I}:\mathfrak{a}_{t})\subset\overline{I}^{\mathfrak{a}^{\left\lceil t\right\rceil}}$.
Further, if $R$ is of prime characteristic $p>0$, then
  1. (c)
    $I^{*\mathfrak{a}^{t}}\subset(I^{*}:\mathfrak{a}_{t})$, and
  2. (d)
    $I^{F\mathfrak{a}^{t}}=\left(I^{F\mathfrak{a}^{t-1}}:\mathfrak{a}\right)$ and ${}^{\mathfrak{a}^{t}}I^{F}=\left({}^{\mathfrak{a}^{t-1}}(\mathfrak{a}I)^{F}:\mathfrak{a}\right)$ if $t>\mu(\mathfrak{a})$.
In Section 4, we carefully study the closures in the setting of affine semigroup rings. In particular, we are able to provide explicit calculations of the adjusted closures in terms of the convex geometry of the associated exponent sets of the ideals involved.
Theorem (Corollary 4.6, Theorem 4.8).
Fix a real number $t>0$ and a field $k$ of prime characteristic $p>0$. Let $\sigma^{\vee}\subseteq\mathbb{R}^{d}$ be a rational polyhedral cone, $S=\sigma^{\vee}\cap\mathbb{Z}^{d}$ the semigroup of lattice points in $\sigma^{\vee}$, $R=k[S]$, and let $I$ and $\mathfrak{a}$ be monomial ideals in $R$. In particular, there are finite subsets $\mathcal{A}$ and $\mathcal{B}$ of $S$ so that $I=(x^{\mathbf{b}}\mid\mathbf{b}\in\mathcal{B})$ and $\mathfrak{a}=(x^{\mathbf{a}}\mid\mathbf{a}\in\mathcal{A})$. We write $H$ for the convex hull of $\mathcal{A}$ in $\mathbb{R}^{d}$. For a monomial $x^{\mathbf{v}}\in R$, we have
  1. (a)
    $x^{\mathbf{v}}\in I^{F\mathfrak{a}^{t}}$ if and only if $x^{\mathbf{v}}\in I^{*\mathfrak{a}^{t}}$ if and only if $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$,
  2. (b)
    $x^{\mathbf{v}}\in{{}^{\mathfrak{a}^{t}}I^{*}}$ if and only if $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$,
  3. (c)
    if $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}I^{F}$ then $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$,
  4. (d)
    if there exists $s>t$ such that $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$, then $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}I^{F}$, and
  5. (e)
    $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$ if and only if $\mathbf{v}+tH\subset\operatorname{Hull}(\operatorname{Exp}I)$.
These theorems show that in the affine semigroup ring case, the formulas which compute the $\mathfrak{a}^{t}$-tight and integral closures are independent of the characteristic of the underlying field. However, we show in an example (see Proposition 4.12) that the $\mathfrak{a}^{t}$-Frobenius closure can depend on the characteristic. We also show in the affine semigroup setting that the joint Hilbert-Kunz multiplicity introduced by Vraciu is computable via the volume of a certain polytope, which reflects the situation for usual Hilbert-Kunz multiplicity by work of K. Eto [Eto02, Theorem 2.2] and K-i. Watanabe [Wat00, Theorem 2.1].
Theorem (Theorem 4.3).
Fix a real number $t>0$ and a field $k$ of prime characteristic $p>0$. Let $S\subset\mathbb{N}^{d}$ be an affine semigroup, $R=k[S]$, and let $I$ and $\mathfrak{a}$ be monomial ideals in $R$ of finite colength. Let $\mathcal{A}$ and $\mathcal{B}$ be finite subsets of $S$ such that $I=(x^{\mathbf{b}}\mid\mathbf{b}\in\mathcal{B})$ and $\mathfrak{a}=(x^{\mathbf{a}}\mid\mathbf{a}\in\mathcal{A})$. We write $H$ for the convex hull of $\mathcal{A}$ and $\sigma^{\vee}$ for the cone defining $S$. If we set
$\mathcal{P}_{t}=\sigma^{\vee}\setminus\left(\mathcal{B}+tH+\sigma^{\vee}\right),$
then the joint Hilbert-Kunz multiplicity of $(\mathfrak{a}^{t},I)$ is given by
$e_{HK}(\mathfrak{a}^{t};I)=\lim\limits_{q\rightarrow\infty}\frac{1}{q^{d}}\ell\left(\frac{R}{\mathfrak{a}^{\lceil tq\rceil}I^{[q]}}\right)=\operatorname{vol}\overline{\mathcal{P}_{t}},$
where $\overline{\mathcal{P}_{t}}$ is the closure of $\mathcal{P}_{t}$ in the usual Euclidean topology on $\mathbb{R}^{d}$ and $\operatorname{vol}$ denotes the usual Euclidean volume.
Finally, in order to demonstrate the theory to the fullest, we provide a thorough analysis of the adjusted closures in a specific affine semigroup ring, see Proposition 4.12. These examples demonstrate a “jumping-number” behavior for the adjusted closures in terms of the parameter $t$. For this example, we also compute the joint Hilbert-Kunz multiplicity as a function of the parameter $t$.
Acknowledgments: The authors are grateful to Rankeya Datta, Mark Johnson, Lance E. Miller, and Janet Vassilev for suggestions and comments. The first author was supported by grant number CEX2021-001142-S (Excelencia Severo Ochoa) funded by MICIU/AEI/10.13039/501100011033 (Ministerio de Ciencia, Innovación y Universidades, Spain).
2. Preliminaries
Conventions.
First, $0\in\mathbb{N}$ and $p$ will always denote a positive prime integer. All rings considered will be commutative with unit and Noetherian. For a ring $R$, we will let $R^{\circ}$ denote the multiplicative set of elements outside of the union of the minimal primes of $R$. That is, if
$S=\bigcup_{\mathfrak{p}\in\operatorname{Min}(R)}\mathfrak{p},$
then $R^{\circ}=R\setminus S$. For an ideal $I\subset R$, $\mu(I)$ denotes the minimal number of generators of $I$. Also, non-positive powers of ideals will be interpreted to mean the entire ring. Finally, for an $\mathbb{N}$-graded ring $R=\bigoplus_{n\geq 0}R_{n}$ and a real number $t$, we will use $R_{<t}$ (similarly $R_{\leq t}$, $R_{\geq t}$, $R_{>t}$) to denote the $R_{0}$-submodule of elements of $R$ of degree less than $t$.
We begin with a corollary to the usual Determinantal Trick for integral closure. While a similar result (without proof) appears in [Vas05, Proposition 1.58] for integrally closed ideals, we provide one for completeness. The ideal $IJ:_{R}J$ described below is called the $J$-basically full closure of $I$. See [ERGV23, Section 4] for a modern treatment.
Theorem 2.1 (The Determinantal Trick).
If $I$ and $J$ are ideals in a ring $R$ with $J\cap R^{\circ}\neq\varnothing$, then $(IJ:_{R}J)\subset\overline{I}$, the integral closure of $I$.
Proof.
Let $x\in(IJ:_{R}J)$, so that $xJ\subset IJ$. Let $\varphi:J\to J$ be given by $\varphi(y)=xy$. By [SH06, Lemma 2.1.8], there exist $r_{i}\in I^{i}$ such that $\varphi^{n}+r_{1}\varphi^{n-1}+\cdots+r_{n}=0$. Let $y\in J\cap R^{\circ}$. Now we have that
$0=(\varphi^{n}+r_{1}\varphi^{n-1}+\cdots+r_{n})(y)=yx^{n}+r_{1}yx^{n-1}+\cdots+r_{n}y=(x^{n}+r_{1}x^{n-1}+\cdots+r_{n})y.$
Let $\mathfrak{p}$ be a minimal prime of $R$. Reducing the above equation modulo $\mathfrak{p}$, we have that, in the domain $R/\mathfrak{p}$,
$(x^{n}+r_{1}x^{n-1}+\cdots+r_{n}+\mathfrak{p})(y+\mathfrak{p})=\mathfrak{p}.$
Since $y+\mathfrak{p}\neq\mathfrak{p}$, it must be the case that $x^{n}+r_{1}x^{n-1}+\cdots+r_{n}\in\mathfrak{p}$. This holds for all minimal primes $\mathfrak{p}$, and therefore we have
$x^{n}+r_{1}x^{n-1}+\cdots+r_{n}\in\bigcap_{\mathfrak{p}\in\mathrm{Spec}\,R}\mathfrak{p}=\sqrt{0}.$
Thus, there exists $m\in\mathbb{N}$ such that
$0=(x^{n}+r_{1}x^{n-1}+\cdots+r_{n})^{m}=x^{mn}+s_{1}x^{mn-1}+\cdots+s_{mn}$
where each $s_{j}$ is a sum of terms of the form $r_{i_{1}}r_{i_{2}}\cdots r_{i_{m}}$ with $i_{1}+i_{2}+\cdots+i_{m}=j$. Hence $s_{j}\in I^{j}$. This shows that $x\in\overline{I}$. ∎
We now review the classical notions of tight and Frobenius closure in prime characteristic.
Definition 2.2.
Let $R$ be a ring of prime characteristic $p>0$ and let $I\subset R$ be an ideal.
  • The $e$-th Frobenius power of $I$, denoted $I^{\left[p^{e}\right]},$ is the ideal of $R$ generated by $\{r^{p^{e}}\mid r\in I\}$.
  • The Frobenius closure of $I$, denoted $I^{F}$, is the ideal of $R$ defined by
    $I^{F}=\left\{x\in R\left|\text{ for some }e\in\mathbb{N},\,x^{p^{e}}\in I^{\left[p^{e}\right]}\right.\right\}.$
  • The tight closure of $I$, denoted $I^{*}$, is the ideal of $R$ defined by
    $I^{*}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{ and all }e\gg 0,\,cx^{p^{e}}\in I^{\left[p^{e}\right]}\right.\right\}.$
For an ideal $I$ in a ring $R$ of prime characteristic $p$, we have that $I\subset I^{F}\subset I^{*}\subset\overline{I}$, but each of these containments are typically strict. However, in a regular ring, $I=I^{*}$ for all $I$ by [HH90, Theorem 4.4], so many singularity types in prime characteristic are defined in terms of whether some or all ideals are closed with respect to these operations. For example, a ring is called weakly $F$-regular if $I^{*}=I$ for all ideals $I$.
The computation of tight and integral closure reduce readily to domains (see [HH90, Proposition 6.25.(a)] and [SH06, Proposition 1.1.5]) as elements of the ring $R$ are in the tight or integral closure of an ideal $I$ if and only if they are in the tight or integral closure of $I+\mathfrak{p}/\mathfrak{p}$ in $R/\mathfrak{p}$ where $\mathfrak{p}$ ranges over all minimal primes of $R$.
It is well-known that the Frobenius closure does not exhibit this property, that is, the inclusion
$I^{F}\subset\left\{x\in R\mid x+\mathfrak{p}\in(I+\mathfrak{p})^{F}/\mathfrak{p}\text{ for any }\mathfrak{p}\in\operatorname{Min}R\right\}$
may be proper; we provide an example below.
Example 2.3.
Let $R=\mathbb{F}_{3}[X,Y,Z]/(Y^{6}-X^{3},X^{2}-Z^{2})$ and let $x,y,$ and $z$ denote the images of $X,Y,$ and $Z$ respectively in $R.$ Let $I=(x)$. One may show that $\operatorname{Min}R=\{(x-z,y^{2}-z),(x+z,y^{2}+z)\}$ and that $z\in(x)^{*}\setminus(x)^{F}.$ Indeed, if $z\in(x)^{F},$ then $z^{3^{e}}\in I^{[3^{e}]}=(x^{3^{e}})$ for some $e.$ Write $13^{e}=2k+$ 1 for some $k.$ Then
$z^{3^{e}}=z^{2k+1}=(z^{2})^{k}z=(x^{2})^{k}z.$
This element belongs to the ideal $(x^{3^{e}})$ only if $2k\geq 3^{e},$ which is a contradiction. Note that modulo either minimal prime, the image of $z$ is a unit times the image of $x$, so
$z+\mathfrak{p}\in((x)+\mathfrak{p})/\mathfrak{p}\subset((x)+\mathfrak{p})^{F}/\mathfrak{p},$
but $z\not\in(x)^{F}$.
Finally, before we begin defining the adjusted closure operations, we remind the reader of the rational powers of an ideal, which are related to the normalization of the Rees algebra of that ideal. The rational powers of an ideal were introduced by Lejuene-Jalabert and Tessier in 1973 (see [LJT08]).
Definition 2.4.
Let $R$ be a ring, $\mathfrak{a}\subset R$ be an ideal, and $t=r/s\in\mathbb{Q}_{\geq 0}$. Then, the $t$-th rational power of $\mathfrak{a}$, denoted $\mathfrak{a}_{t}$, is defined by $\mathfrak{a}_{t}=\left\{x\in R\left|\,x^{s}\in\overline{\mathfrak{a}^{r}}\right.\right\}.$
Despite being introduced many years ago, the rational powers of ideals have received little attention until recently. See [SH06, Section 10.5] for a treatment of rational powers. For some modern results utilizing them, see [Ciu20] and [BDHM24].
3. The closure operations
In this section we define four new closure operations whose origin is a formal adaptation of the tight closure adjustments of Hara-Yoshida [HY03] and Vraciu [Vra08] (see 3.9) to integral and Frobenius closures. We also note that in this section we use the word “closure” loosely, in that the Hara-Yoshida versions of the closures are not generally idempotent operations, so that the Hara-Yoshida “closures” may not be closed.
3.1. The $\mathfrak{a}^{t}$-integral closures
In this subsection, $R$ is a ring of arbitrary characteristic, and we will assume the ideals we denote $I$ and $\mathfrak{a}$ intersect $R^{\circ}$ non-trivially. We begin with the adjustments to the ordinary integral closure of an ideal $I$.
Definition 3.1.
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$. Then, the Hara-Yoshida $\mathfrak{a}^{t}$-Integral closure of $I$, denoted $\overline{I}^{\mathfrak{a}^{t}}$, is defined by
$\overline{I}^{\mathfrak{a}^{t}}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{ and all }n\gg 0,\,cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}\right.\right\}.$
Similarly, the Vraciu $\mathfrak{a}^{t}$-Integral closure of $I$, denoted ${}^{\mathfrak{a}^{t}}\overline{I}$, is defined by
${}^{\mathfrak{a}^{t}}\overline{I}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{ and all }n\gg 0,\,cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset\mathfrak{a}^{\lceil tn\rceil}I^{n}\right.\right\}.$
We first note that the “for all $n\gg 0$” in the definitions above may be strengthened to “for all $n$”, so long as $I$ is of positive height.
Proposition 3.2.
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$. Then $x\in\overline{I}^{\mathfrak{a}^{t}}$ (respectively $x\in{}^{\mathfrak{a}^{t}}\overline{I}$) if and only if there exists $c\in R^{\circ}$ such that, for all $n\geq 0$, $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ (respectively $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}\mathfrak{a}^{\lceil tn\rceil}$).
Proof.
The proofs of both parts are similar, so we provide the proof for the Hara-Yoshida version only. Let $x\in\overline{I}^{\mathfrak{a}^{t}}.$ Then there exist $n_{0}\in\mathbb{N}$ and $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for all $n\geq n_{0}.$ Choose $b\in R^{\circ}\cap I^{n_{0}}$. Then $bcx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for all $n\geq 0.$
We now record several basic properties of these closure operations. Notably, as long as $I$ and $\mathfrak{a}$ avoid the minimal primes of $R$, the Vraciu $\mathfrak{a}^{t}$-integral closure agrees with the usual integral closure, but the Hara-Yoshida version is generally larger.
Proposition 3.3.
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$.
  1. (a)
    ${}^{\mathfrak{a}^{t}}\overline{I}$ and $\overline{I}^{\mathfrak{a}^{t}}$ are ideals of $R.$
  2. (b)
    $\overline{I}={}^{\mathfrak{a}^{t}}\overline{I}\subset\overline{I}^{\mathfrak{a}^{t}}.$
  3. (c)
    If $I\subset J$, $\mathfrak{b}\subset\mathfrak{a}$, and $t\leq s$, then $\overline{I}^{\mathfrak{a}^{t}}\subset\overline{J}^{\mathfrak{b}^{s}}$ and ${}^{\mathfrak{a}^{t}}\overline{I}\subset{}^{\mathfrak{a}^{s}}\overline{J}.$
  4. (d)
    $I^{*\mathfrak{a}^{t}}\subset\overline{I}^{\mathfrak{a}^{t}}$ and ${}^{\mathfrak{a}^{t}}I^{*}\subset{}^{\mathfrak{a}^{t}}\overline{I}.$
Proof.
We will prove each in turn.
  1. (a)
    In order to show that $\overline{I}^{\mathfrak{a}^{t}}$ is an ideal, it is enough to check that if $x,y\in\overline{I}^{\mathfrak{a}^{t}},$ then $x+y\in\overline{I}^{\mathfrak{a}^{t}}.$ As $x,y\in\overline{I}^{\mathfrak{a}^{t}},$ there exist $c,d\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ and $dy^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for all $n\geq 0.$ Choose $c^{\prime}\in\mathfrak{a}^{2}\cap R^{\circ}.$ Then
    $c^{\prime}cd(x+y)^{n}\mathfrak{a}^{\lceil tn\rceil}=\sum_{j=0}^{n}\binom{n}{j}(cx^{j})(dy^{n-j})c^{\prime}\mathfrak{a}^{\lceil tn\rceil}$ $\subset\sum_{j=0}^{n}\binom{n}{j}(cx^{j})(dy^{n-j})\mathfrak{a}^{\lceil tn\rceil+2}$
    $\subset\sum_{j=0}^{n}\binom{n}{j}(cx^{j}\mathfrak{a}^{\lceil tj\rceil})(dy^{n-j}\mathfrak{a}^{\lceil t(n-j)\rceil})\subset I^{n}.$
    Thus, $\overline{I}^{\mathfrak{a}^{t}}$ is an ideal. Now let $x,y\in{}^{\mathfrak{a}^{t}}\overline{I}.$ Then there exist $c,d\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset\mathfrak{a}^{\lceil tn\rceil}I^{n},$ and $dy^{n}\mathfrak{a}^{\lceil tn\rceil}\subset\mathfrak{a}^{\lceil tn\rceil}I^{n}$ for all $n\geq 0.$ Choose $c^{\prime}\in\mathfrak{a}^{2}\cap R^{\circ}.$ Then
    $c^{\prime}cd(x+y)^{n}\mathfrak{a}^{\lceil tn\rceil}=\sum_{j=0}^{n}\binom{n}{j}(cx^{j})(dy^{n-j})c^{\prime}\mathfrak{a}^{\lceil tn\rceil}$ $\subset\sum_{j=0}^{n}\binom{n}{j}(cx^{j}\mathfrak{a}^{\lceil tj\rceil})(dy^{n-j}\mathfrak{a}^{\lceil t(n-j)\rceil})$
    $\subset\sum_{j=0}^{n}\mathfrak{a}^{\lceil tj\rceil}\mathfrak{a}^{\lceil t(n-j)\rceil}I^{n}\subset\mathfrak{a}^{\lceil tn\rceil}I^{n}.$
    Thus, ${}^{\mathfrak{a}^{t}}\overline{I}$ is an ideal.
  2. (b)
    Since $I^{n}\subset I^{n}\mathfrak{a}^{\lceil tn\rceil}:\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}:\mathfrak{a}^{\lceil tn\rceil}$, for all $n$, we have $\overline{I}\subset{}^{\mathfrak{a}^{t}}\overline{I}\subset\overline{I}^{\mathfrak{a}^{t}}$. Now let $x\in{}^{\mathfrak{a}^{t}}\overline{I}.$ Then there exists $c\in R^{\circ}$ such that for all $n$, $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset\mathfrak{a}^{\lceil tn\rceil}I^{n}$. Using Theorem 2.1, we get $cx^{n}\in\overline{I^{n}}$ for all $n$ implying that $x\in\overline{I}$ (see [SH06, Corollary 6.8.12]).
  3. (c)
    Since $I\subset J$, $I^{n}\subset J^{n}$ for all $n$, so we have $I^{n}:\mathfrak{a}^{\lceil tn\rceil}\subset J^{n}:\mathfrak{b}^{\lceil sn\rceil}$ for all $n$. Thus, $\overline{I}^{\mathfrak{a}^{t}}\subset\overline{J}^{\mathfrak{b}^{s}}$. Similarly, $I^{n}\mathfrak{a}^{\lceil tn\rceil}:\mathfrak{a}^{\lceil tn\rceil}\subset J^{n}\mathfrak{a}^{\lceil tn\rceil}:\mathfrak{a}^{\lceil tn\rceil}\subset J^{n}\mathfrak{a}^{\lceil sn\rceil}:\mathfrak{a}^{\lceil sn\rceil}$ for all $n$, so ${}^{\mathfrak{a}^{t}}\overline{I}\subset{}^{\mathfrak{a}^{s}}\overline{J}$.
  4. (d)
    Let $x\in I^{*\mathfrak{a}^{t}}.$ Then there exists $c\in R^{\circ}$ such that for all $e\gg 0,$ $cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\subset I^{p^{e}}$ implying that $x\in\overline{I}^{\mathfrak{a}^{t}}.$ One can show the other containment similarly. ∎
Remark 3.4.
Due to the fact that in most cases of consideration (e.g. nonzero ideals in an integral domain) we have that ${}^{\mathfrak{a}^{t}}\overline{I}=\overline{I}$, we will not typically have reason to discuss the Vraciu $\mathfrak{a}^{t}$-integral closure later in the article, and will refer to $\overline{I}^{\mathfrak{a}^{t}}$ as simply the $\mathfrak{a}^{t}$-integral closure of $I$.
Proposition 3.5.
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$.
  1. (a)
    If $t\geq 1$ then $\overline{I}^{\mathfrak{a}^{t}}\subset\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)$, and we have equality if $\mathfrak{a}$ is a principal ideal or $t\in\mathbb{Z}$.
  2. (b)
    If $t\in\mathbb{Z}_{\geq 1}$, then $\overline{I}^{\mathfrak{a}^{t}}=(\overline{I}:\mathfrak{a}^{t}).$
  3. (c)
    If $t\in\mathbb{Q}_{>0}$, then $\overline{I}^{\mathfrak{a}^{t}}\subset(\overline{I}:\mathfrak{a}_{t})\subset\overline{I}^{\mathfrak{a}^{\lceil t\rceil}}.$
Proof.
We will show each in turn.
  1. (a)
    Let $x\in\overline{I}^{\mathfrak{a}^{t}}$, so that there exists $c\in R^{\circ}$ such that for all $n\gg 0$, $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$. For any $y\in\mathfrak{a}$ and $n$ we have
    $c(xy)^{n}\mathfrak{a}^{\lceil(t-1)n\rceil}\subset cx^{n}\mathfrak{a}^{n}\mathfrak{a}^{\lceil(t-1)n\rceil}=cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}.$
    Therefore $xy\in\overline{I}^{\mathfrak{a}^{t}}$. Hence $x\in\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)$.
    Suppose $\mathfrak{a}=(f)$ is principal, and let $x\in\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)$, so that $xf\in\overline{I}^{\mathfrak{a}^{t-1}}$. Let $c\in R^{\circ}$ such that $c(xf)^{n}\mathfrak{a}^{\lceil(t-1)n\rceil}\subset I^{n}$ for $n$. Then for all $n$,
    $cx^{n}\mathfrak{a}^{\lceil tn\rceil}=cx^{n}(f^{\lceil tn\rceil})=c(xf)^{n}(f^{\lceil(t-1)n\rceil})=c(xf)^{n}\mathfrak{a}^{\lceil(t-1)n\rceil}\subset I^{n}.$
    Therefore $x\in\overline{I}^{\mathfrak{a}^{t}}$, and so $\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)\subset\overline{I}^{\mathfrak{a}^{t}}$, and we have equality.
    Finally, suppose $t\in\mathbb{Z}$ and $x\in\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)$. Let $y_{1},\ldots,y_{r}$ be the generators of $\mathfrak{a}$. For each $i$, $xy_{i}\in\overline{I}^{\mathfrak{a}^{t-1}}$, and so there exists $c_{i}\in R^{\circ}$ such that for all $n\geq 0$, $c_{i}(xy)^{n}\mathfrak{a}^{(t-1)n}\subset I^{n}$. Let $c=\prod_{i}c_{i}\in R^{\circ}$, and fix $n\geq 0$. All generators of $\mathfrak{a}^{n}$ are of the form $\prod_{i}y_{i}^{n_{i}}$ with $\sum_{i}n_{i}=n$. For such a generator, we have that
    $cx^{n}\left(\prod_{i}y_{i}^{n_{i}}\right)\mathfrak{a}^{(t-1)n}=\prod_{i}\left(c_{i}x^{n_{i}}y^{n_{i}}\mathfrak{a}^{(t-1)n_{i}}\right)\subset\prod_{i}I^{n_{i}}=I^{n}.$
    Therefore $cx^{n}\mathfrak{a}^{tn}=cx^{n}\mathfrak{a}^{n}\mathfrak{a}^{(t-1)n}\subset I^{n}$, meaning $x\in\overline{I}^{\mathfrak{a}^{t}}$. Hence $\left(\overline{I}^{\mathfrak{a}^{t-1}}:\mathfrak{a}\right)\subset\overline{I}^{\mathfrak{a}^{t}}$, and we have equality.
  2. (b)
    Repeatedly apply the previous part $t$ times.
  3. (c)
    Suppose $x$ is such that, for some $c\in R^{\circ}$ and all $n\geq 0$, we have $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$, and let $y\in\mathfrak{a}_{t}$. By [Tay21, Lemma 2.8], there exists $d\in R^{\circ}$ such that for all $n\gg 0$, $dy^{n}\in\mathfrak{a}^{\lceil tn\rceil}$. Now $cd\in R^{\circ}$ and for $n\gg 0$, $cd(xy)^{n}=cx^{n}dy^{n}\in cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$. Therefore $xy\in\overline{I}$ and so $\overline{I}^{\mathfrak{a}^{t}}\subset(\overline{I}:\mathfrak{a}_{t})$.
    Suppose that $x\in(\overline{I}:\mathfrak{a}_{t})$. Let $c\in R^{\circ}$ such that for $n\geq 0$, $c(\overline{I})^{n}\subset I^{n}$. Now, for $n\gg 0$,
    $cx^{n}\mathfrak{a}^{\lceil t\rceil n}=c(x\mathfrak{a}^{\lceil t\rceil})^{n}\subset c(x\mathfrak{a}_{\lceil t\rceil})^{n}\subset c(x\mathfrak{a}_{t})^{n}\subset c(\overline{I})^{n}=I^{n}.$
    Therefore $x\in\overline{I}^{\mathfrak{a}^{\lceil t\rceil}}$. ∎
Remark 3.6.
The reverse containment $(\overline{I}:\mathfrak{a}_{t})\subset\overline{I}^{\mathfrak{a}^{t}}$ in 3.5 is not true in general if $t$ is not an integer. For example, let $R=k[x,y]$, $\mathfrak{a}=(x,y)$, $I=(x^{2},y^{2})$, and $21<t<$ 2. Here $\mathfrak{a}_{t}=(x^{2},xy,y^{2})=\overline{I}$, and so $(\overline{I}:\mathfrak{a}_{t})=R$. However, $1\notin\overline{I}^{\mathfrak{a}^{t}}$. If it were, then there would exist a nonzero $c\in R$ such that for all large $n$, $c\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$. However, for any $n\in\mathbb{N}$, $\mathfrak{a}^{\lceil tn\rceil}$ is generated by elements of degree $\lceil tn\rceil$, and so $c\mathfrak{a}^{\lceil tn\rceil}$ contains elements of degree $\deg(c)+\lceil tn\rceil$. Since $I$ is generated by elements of degree 2, $I^{n}$ is generated by homogeneous elements of degree $2n$. Thus, we would have that $2n\leq\deg(c)+\lceil tn\rceil$ for all $n$, meaning that $\deg(c)\geq\lceil(2-t)n\rceil$ for all $n$, a contradiction. Therefore $(\overline{I}:\mathfrak{a}_{t})\not\subset\overline{I}^{\mathfrak{a}^{t}}$.
We may use 3.5 to observe that the Hara-Yoshida integral closure is not an idempotent operation on ideals in general.
Example 3.7.
Let $R$ be a ring, $I\subset R$ an ideal and $f\in R$ a nonzerodivisor. By 3.5, $\overline{I}^{(f)}=(\overline{I}:f)$ and
$\overline{\Big{(}\overline{I}^{(f)}\Big{)}}^{(f)}=\Big{(}\overline{I}^{(f)}:f\Big{)}=((\overline{I}:f):f)=(\overline{I}:f^{2}).$
Since in general $(\overline{I}:f)$ may be strictly smaller than $(\overline{I}:f^{2})$, we have that $\overline{I}^{\mathfrak{a}^{t}}$ may be strictly smaller than $\overline{\left(\overline{I}^{\mathfrak{a}^{t}}\right)}^{\mathfrak{a}^{t}}.$
We conclude this subsection by demonstrating that computation of the $\mathfrak{a}^{t}$-integral closure may be reduced to computation in domains via modding out by minimal primes, unlike Frobenius closure (see 2.3).
Proposition 3.8.
Let $\mathfrak{a},I$ be ideals in a reduced ring $R$ and let $t\in\mathbb{R}_{>0}$. An element $x\in\overline{I}^{\mathfrak{a}^{t}}$ if and only if the image of $x$ in $R/\mathfrak{p}$ is in $\overline{((I+\mathfrak{p})/\mathfrak{p})}^{\mathfrak{b}^{t}}$ for all minimal primes $\mathfrak{p}$ of $R.$ Here, $\mathfrak{b}$ denotes the image of $\mathfrak{a}$ in $R/\mathfrak{p}.$
Proof.
Let $x\in\overline{I}^{\mathfrak{a}^{t}}.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for all $n.$ Thus, for any minimal prime $\mathfrak{p}$ of $R,$ the image of $c$ in $R/\mathfrak{p}$ is in $(R/\mathfrak{p})^{\circ}$ and $cx^{n}\mathfrak{a}^{\lceil tn\rceil}+\mathfrak{p}\subset I^{n}+\mathfrak{p}$ for all $n$ implying that the image of $x$ in $R/\mathfrak{p}$ is in $\overline{((I+\mathfrak{p})/\mathfrak{p})}^{\mathfrak{b}^{t}}.$
Conversely, assume that for each minimal prime $\mathfrak{p}$ of $R,$ image of $x$ in $R/\mathfrak{p}$ is in $\overline{((I+\mathfrak{p})/\mathfrak{p})}^{\mathfrak{b}^{t}}.$ Let $c_{\mathfrak{p}}\in R\setminus\mathfrak{p}$ such that $c_{\mathfrak{p}}x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset c_{\mathfrak{p}}x^{n}\mathfrak{a}^{\lceil tn\rceil}+\mathfrak{p}\subset I^{n}+\mathfrak{p}$ for all $n.$ By prime avoidance, there exists a non-zero element $d_{\mathfrak{p}}\in R$ which is not in $\mathfrak{p}$ but is in every other minimal prime of $R.$ Then $d_{\mathfrak{p}}c_{\mathfrak{p}}x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}+\sqrt{0},$ for all $n.$ Set $c^{\prime}=\sum_{\mathfrak{p}\in\operatorname{MinR}}d_{\mathfrak{p}}c_{\mathfrak{p}}.$ Then $c^{\prime}\in R^{\circ}$ and for all $n$, we have
$c^{\prime}x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}+\sqrt{0}=I^{n}$
implying that $x\in\overline{I}^{\mathfrak{a}^{t}}.$
3.2. The $\mathfrak{a}^{t}$-tight closures
Throughout this subsection, we assume $R$ is of prime characteristic $p$. Recall the following definitions from [HY03] and [Vra08].
Definition 3.9.
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$. Then, the Hara-Yoshida $\mathfrak{a}^{t}$-tight closure of $I$, denoted $I^{*\mathfrak{a}^{t}},$ is defined by
$I^{*\mathfrak{a}^{t}}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{and all }e\gg 0,\,cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{\left[p^{e}\right]}\right.\right\}.$
Similarly, the Vraciu $\mathfrak{a}^{t}$-tight closure of $I$, denoted ${}^{\mathfrak{a}^{t}}I^{*}$, is defined by
${}^{\mathfrak{a}^{t}}I^{*}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{ and all }e\gg 0,\,cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset\mathfrak{a}^{\lceil tp^{e}\rceil}I^{\left[p^{e}\right]}\right.\right\}.$
We note that like the Hara-Yoshida integral closure, the Hara-Yoshida $\mathfrak{a}^{t}$-tight closure is related to rational powers of $\mathfrak{a}$; compare the following result with 3.5.
Proposition 3.10.
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{Q}_{>0}$. Then $I^{*\mathfrak{a}^{t}}\subset(I^{*}:\mathfrak{a}_{t})$.
Proof.
Let $x\in I^{*\mathfrak{a}^{t}}$ with $c\in R^{\circ}$ such that $cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}$ for all $e\gg 0$, and let $y\in\mathfrak{a}_{t}$. By [Tay21, Lemma 2.8], there exists $d\in R^{\circ}$ such that for all $e\gg 0$, $dy^{p^{e}}\in\mathfrak{a}^{\lceil tp^{e}\rceil}$. As $cd(xy)^{p^{e}}\in cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}$ for all $e$ large, we get $xy\in I^{*}$ and so $I^{*\mathfrak{a}^{t}}\subset(I^{*}:\mathfrak{a}_{t})$. ∎
The following result shows that, like tight closure, the computation of $\mathfrak{a}^{t}$-tight closure reduces to computation in domains via modding out by minimal primes. Note that while we mention the result for the $t=1$ case, it holds similarly for $t\neq 1$ as well.
Proposition 3.11.
Let $\mathfrak{a},I\subset R$ be ideals with $\mathfrak{a}\cap R^{\circ}\neq\varnothing.$ Then, $x\in$ ${}^{\mathfrak{a}}I^{*}$ ($x\in I^{*\mathfrak{a}}$) if and only if, for each minimal prime $\mathfrak{p}$ of $R$, we have $x+\mathfrak{p}$ in ((${}^{\mathfrak{b}}((I+\mathfrak{p})/\mathfrak{p})^{*}$ (or $((I+\mathfrak{p})/\mathfrak{p})^{*\mathfrak{b}}$), where $\mathfrak{b}$ denotes the image of $\mathfrak{a}$ in $R/\mathfrak{p}.$
Proof.
Let $x\in$ ${}^{\mathfrak{a}}I^{*}.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{p^{e}}\mathfrak{a}^{p^{e}}\subset\mathfrak{a}^{p^{e}}I^{[p^{e}]}$ for all large $e.$ Thus, for any minimal prime $\mathfrak{p}$ of $R,$ image of $c$ in $R/\mathfrak{p}$ is in $(R/\mathfrak{p})^{\circ}$ and $cx^{p^{e}}\mathfrak{a}^{p^{e}}+\mathfrak{p}\subset\mathfrak{a}^{p^{e}}I^{[p^{e}]}+\mathfrak{p}$ for all large $e$ implying that the image of $x$ in $R/\mathfrak{p}$ is in ((${}^{\mathfrak{b}}((I+\mathfrak{p})/\mathfrak{p})^{*}.$
Conversely, assume that for each minimal prime $\mathfrak{p}$ of $R,$ image of $x$ in $R/\mathfrak{p}$ is in ((${}^{\mathfrak{b}}((I+\mathfrak{p})/\mathfrak{p})^{*}.$ Let $c_{\mathfrak{p}}\in R\setminus\mathfrak{p}$ such that $c_{\mathfrak{p}}x^{p^{e}}\mathfrak{a}^{p^{e}}\subset c_{\mathfrak{p}}x^{p^{e}}\mathfrak{a}^{p^{e}}+\mathfrak{p}\subset\mathfrak{a}^{p^{e}}I^{[p^{e}]}+\mathfrak{p}$ for all $e$ large. By Prime Avoidance, there exists a non-zero element $d_{\mathfrak{p}}\in R$ which is not in $\mathfrak{p}$ but is in every other minimal prime of $R.$ Then $d_{\mathfrak{p}}c_{\mathfrak{p}}x^{p^{e}}\mathfrak{a}^{p^{e}}\subset\mathfrak{a}^{p^{e}}I^{[p^{e}]}+\sqrt{0},$ for all $e$ large. Set $c^{\prime}=\sum_{\mathfrak{p}}d_{\mathfrak{p}}c_{\mathfrak{p}}.$ Then $c^{\prime}\in R^{\circ}$ and for all $e$ large,
$c^{\prime}x^{p^{e}}\mathfrak{a}^{p^{e}}\subset\mathfrak{a}^{p^{e}}I^{[p^{e}]}+\sqrt{0}.$
Since $R$ is a Noetherian ring, there exists $e^{\prime}$ such that $(\sqrt{0})^{[p^{e^{\prime}}]}=(0).$ Thus, for all $e$ large, we have $c^{\prime\prime}x^{p^{e+e^{\prime}}}(\mathfrak{a}^{p^{e}})^{[p^{e^{\prime}}]}\subset(\mathfrak{a}^{p^{e}})^{[p^{e^{\prime}}]}I^{[p^{e+e^{\prime}}]}$. From [Tay21, Lemma 2.1], we know that $\mathfrak{a}^{p^{e+e^{\prime}}}\mathfrak{a}^{(\mu-1)p^{e^{\prime}}}\subset(\mathfrak{a}^{[p^{e^{\prime}}]})^{p^{e}},$ where $\mu$ denotes the minimal number of generators of $\mathfrak{a}.$ Hence, for any $b\in\mathfrak{a}^{(\mu-1)p^{e^{\prime}}}\cap R^{\circ},$ and for all $e$ large,
$bc^{\prime\prime}x^{p^{e+e^{\prime}}}\mathfrak{a}^{p^{e+e^{\prime}}}\subset c^{\prime\prime}x^{p^{e+e^{\prime}}}(\mathfrak{a}^{[p^{e^{\prime}}]})^{p^{e}}\subset(\mathfrak{a}^{p^{e}})^{[p^{e^{\prime}}]}I^{[p^{e+e^{\prime}}]}\subset\mathfrak{a}^{p^{e+e^{\prime}}}I^{[p^{e+e^{\prime}}]}.$
As $bc^{\prime\prime}\in R^{\circ},$ we get $x\in$ ${}^{\mathfrak{a}}I^{*}$. The proof for the ideal $I^{*\mathfrak{a}}$ follows similarly. ∎
3.3. The $\mathfrak{a}^{t}$-Frobenius closures
Throughout this subsection, we assume $R$ is of prime characteristic $p$.
Definition 3.12.
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$. Then, the Hara-Yoshida $\mathfrak{a}^{t}$-Frobenius closure of $I$, denoted $I^{F\mathfrak{a}^{t}},$ is defined by
$I^{F\mathfrak{a}^{t}}=\left\{x\in R\left|\text{ for all }e\gg 0,\,x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{\left[p^{e}\right]}\right.\right\}.$
Similarly, the Vraciu $\mathfrak{a}^{t}$-Frobenius closure of $I$, denoted ${}^{\mathfrak{a}^{t}}I^{F}$, is defined by
${}^{\mathfrak{a}^{t}}I^{F}=\left\{x\in R\left|\text{ for all }e\gg 0,\,x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset\mathfrak{a}^{\lceil tp^{e}\rceil}I^{\left[p^{e}\right]}\right.\right\}.$
We now record several basic properties of these operations.
Proposition 3.13.
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$. Then, we have the following.
  1. (a)
    ${}^{\mathfrak{a}^{t}}I^{F}$ and $I^{F\mathfrak{a}^{t}}$ are ideals of $R.$
  2. (b)
    $I^{F}\subset{}^{\mathfrak{a}^{t}}I^{F}\subset I^{F\mathfrak{a}^{t}}.$
  3. (c)
    If $I\subset J$, $\mathfrak{b}\subset\mathfrak{a}$, and $t\leq s$, then $I^{F\mathfrak{a}^{t}}\subset J^{F\mathfrak{b}^{s}}$ and ${}^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{s}}J^{F}.$
  4. (d)
    (${}^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{F})^{F}={}^{\mathfrak{a}^{t}}I^{F}.$
Proof.
We will prove each in turn.
  1. (a)
    Notice that $x\in I^{F\mathfrak{a}^{t}}$ if and only if $x^{p^{e}}\in I^{\left[p^{e}\right]}:\mathfrak{a}^{\lceil tp^{e}\rceil}$ for all $e\gg 0$, from whence it is obvious that $I^{F\mathfrak{a}^{t}}$ is an ideal of $R$. Similarly, $x\in{}^{\mathfrak{a}^{t}}I^{F}$ if and only if $x^{p^{e}}\in I^{\left[p^{e}\right]}\mathfrak{a}^{\lceil tp^{e}\rceil}:\mathfrak{a}^{\lceil tp^{e}\rceil}$, so ${}^{\mathfrak{a}^{t}}I^{F}$ is also an ideal of $R$.
  2. (b)
    Since $I^{\left[p^{e}\right]}\subset I^{\left[p^{e}\right]}\mathfrak{a}^{\lceil tp^{e}\rceil}:\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{\left[p^{e}\right]}:\mathfrak{a}^{\lceil tp^{e}\rceil}$, for all $e\in\mathbb{N}$, we have $I\subset I^{F}\subset{}^{\mathfrak{a}^{t}}I^{F}\subset I^{F\mathfrak{a}^{t}}$.
  3. (c)
    Since $I\subset J$, $I^{\left[p^{e}\right]}\subset J^{\left[p^{e}\right]}$ for all $e$, we have $I^{\left[p^{e}\right]}:\mathfrak{a}^{\lceil tp^{e}\rceil}\subset J^{\left[p^{e}\right]}:\mathfrak{b}^{\lceil sp^{e}\rceil}$ for all $e$. Thus, $I^{F\mathfrak{a}^{t}}\subset J^{F\mathfrak{b}^{s}}$. Similarly, $I^{\left[p^{e}\right]}\mathfrak{a}^{\lceil tp^{e}\rceil}:\mathfrak{a}^{\lceil tp^{e}\rceil}\subset J^{\left[p^{e}\right]}\mathfrak{a}^{\lceil tp^{e}\rceil}:\mathfrak{a}^{\lceil tp^{e}\rceil}\subset J^{[p^{e}]}\mathfrak{a}^{\lceil sp^{e}\rceil}:\mathfrak{a}^{\lceil sp^{e}\rceil}$ for all $e$, implying that ${}^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{s}}J^{F}$.
  4. (d)
    From part b), it follows that ${}^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{F})^{F}.$ Now, let $x\in{}^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{F})^{F}.$ Then for all $e\gg 0,$ $x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset\mathfrak{a}^{\lceil tp^{e}\rceil}({}^{\mathfrak{a}^{t}}I^{F})^{[p^{e}]}\subset\mathfrak{a}^{\lceil tp^{e}\rceil}I^{[p^{e}]},$ where the last containment holds by checking the generators of ${}^{\mathfrak{a}^{t}}I^{F}$ and using that ${}^{\mathfrak{a}^{t}}I^{F}$ is a finitely generated $R$-module. ∎
Note the differences in the montonicity properties of the Hara-Yoshida and Vraciu $\mathfrak{a}^{t}$-Frobenius closures outlined in (c) above. In fact, as for the Vraciu $\mathfrak{a}^{t}$-tight closure (see [Vra08, Observation 5.1]), the following example illustrates that it is not always true that ${}^{\mathfrak{a}}I^{F}\subset$ ${}^{\mathfrak{b}}I^{F}$ if $\mathfrak{b}\subset\mathfrak{a}.$
Example 3.14.
Let $R=k[x,y],$ where $k$ is a field of prime characteristic $p>0.$ Let $\mathfrak{a}=(x,y)^{2},$ $\mathfrak{b}=(f),$ where $f\in\mathfrak{a}$ and $I=(x^{2},y^{2}).$ From Proposition 3.16(b), we know that ${}^{\mathfrak{b}}I^{F}=I^{F}=I$ as $R$ is a regular ring. But $xy\in$ ${}^{\mathfrak{a}}I^{F}$ as $(xy)^{p^{e}}\mathfrak{a}^{p^{e}}\subset I^{[p^{e}]}\mathfrak{a}^{p^{e}}$ for all $e$.
Unlike $\mathfrak{a}^{t}$-integral closure and $\mathfrak{a}^{t}$-tight closure, $\mathfrak{a}^{t}$-Frobenius closure can’t be computed using the corresponding closure ideal after going modulo the minimal primes. We use 2.3 to check that 3.11 may not hold true for ${}^{\mathfrak{a}^{t}}I^{F}$ and $I^{F\mathfrak{a}^{t}}$.
Example 3.15.
Let $R=\mathbb{F}_{3}[X,Y,Z]/(Y^{6}-X^{3},X^{2}-Z^{2})$ and let $x,y,$ and $z$ denote the images of $X,Y,$ and $Z$ respectively in $R.$ Let $I=(x)$ and $\mathfrak{a}=(x,y^{2})$. We claim that $z\notin$ ${}^{\mathfrak{a}}I^{F}.$ Indeed, if $z\in$ ${}^{\mathfrak{a}}I^{F},$ then $z^{3^{e}}(x,y^{2})^{3^{e}}\subset\mathfrak{a}^{3^{e}}I^{[3^{e}]}$ for some $e.$ In particular, $z^{3^{e}}x^{3^{e}}\in\mathfrak{a}^{3^{e}}I^{[3^{e}]}=(x,y^{2})^{3^{e}}(x^{3^{e}})$ for some $e.$ Write $13^{e}=2k+$ 1 for some $k.$ Then
$z^{3^{e}}x^{3^{e}}=z^{2k+1}x^{3^{e}}=z(z^{2})^{k}x^{3^{e}}=zx^{2k}x^{3^{e}}=zx^{3^{e}-1}x^{3^{e}}.$
Since $x$ is a nonzerodivisor, we get $z\in(x,y^{2})^{3^{e}}:x^{3^{e}-1}=(x,y^{2})$ which is not true. Thus, the claim holds. Note that $\operatorname{Min}R=\{(x-z,y^{2}-z),(x+z,y^{2}+z)\}$ and so modulo either minimal prime, the image of $z$ is the same as the image of $x.$ Thus, $z+\mathfrak{p}\in((x)+\mathfrak{p})/\mathfrak{p}\subset$ ((${}^{\mathfrak{a}}((x)+\mathfrak{p})^{F}/\mathfrak{p}$, but $z\not\in$ (${}^{\mathfrak{a}}(x)^{F}$.
Now let $I=(x+y^{2})$ and $\mathfrak{a}=(x,y^{2}).$ Then one can check that modulo any minimal prime $\mathfrak{p},$ the element $(1+\mathfrak{p})/\mathfrak{p}$ is in $((I+\mathfrak{p})/\mathfrak{p})^{F\mathfrak{a}}$ but not in $I^{F\mathfrak{a}}.$ In particular, it is enough to show that $(\mathfrak{a}^{3^{e}}+\mathfrak{p})/\mathfrak{p}\subset(I^{[3^{e}]}+\mathfrak{p})/\mathfrak{p}$ for every minimal prime $\mathfrak{p}$ and for all $e$ large but $\mathfrak{a}^{3^{e}}\not\subset I^{[3^{e}]}$ for any $e.$ Indeed, modulo any minimal prime $\mathfrak{p},$
$\frac{I^{[3^{e}]}+\mathfrak{p}}{\mathfrak{p}}=\frac{(x^{3^{e}}+(y^{2})^{3^{e}})+\mathfrak{p}}{\mathfrak{p}}=\frac{(x^{3^{e}}+x^{3^{e}})+\mathfrak{p}}{\mathfrak{p}}=\frac{(x^{3^{e}})+\mathfrak{p}}{\mathfrak{p}}=\frac{\mathfrak{a}^{3^{e}}+\mathfrak{p}}{\mathfrak{p}},$
for all $e.$ Whereas in the ring $R$, if $\mathfrak{a}^{3^{e}}\subset I^{[3^{e}]}$ for some $e,$ then $x^{3^{e}-1}y^{2}\in(x^{3^{e}}+(y^{2})^{3^{e}})=(x^{3^{e}}).$ But as $x$ is a nonzerodivisor, this means, $y^{2}\in(x)$ which is not true.
As for the $\mathfrak{a}^{t}$-tight and integral closures, if $\mathfrak{a}=R$, then $I^{FR^{t}}=\,^{R^{t}}I^{F}=I^{F}$.
Proposition 3.16.
The following observations demonstrate several cases where the adjusted Frobenius closures are simple to calculate or trivially related to another operation.
  1. (a)
    We have $I^{F\mathfrak{a}}\subset(I^{F}:\mathfrak{a}),$ and the inclusion is an equality if $\mathfrak{a}$ is a principal ideal. In particular, if $\mathfrak{a}=(f)$, we have $(I^{F\mathfrak{a}})^{F\mathfrak{a}}\neq I^{F\mathfrak{a}}$ when $(R,\mathfrak{m})$ is local, $I$ is $\mathfrak{m}$-primary, and $f\in\mathfrak{m}\setminus I^{F}.$
  2. (b)
    If $\mathfrak{a}=(f)$ is a principal ideal, and $f$ is a non-zerodivisor on $R$, then $\,{}^{\mathfrak{a}^{t}}I^{F}=I^{F}.$
  3. (c)
    Let $t>\mu(\mathfrak{a})$ a real number, where $\mu(\mathfrak{a})$ denotes the minimal number of generators of $\mathfrak{a}$. Then
    1. (i)
      $I^{F\mathfrak{a}^{t}}=(I^{F\mathfrak{a}^{t-1}}:\mathfrak{a}),$ and
    2. (i)
      ${}^{\mathfrak{a}^{t}}I^{F}=(^{\mathfrak{a}^{t-1}}(\mathfrak{a}I)^{F}:\mathfrak{a}).$
  4. (d)
    If $\mathfrak{a}\subset\sqrt{0}$, then for all ideals $I\subset R$ and $t\in\mathbb{R}_{>0}$, we have ${}^{\mathfrak{a}^{t}}I^{F}=R$.
Proof.
  1. (a)
    The first statement holds using the definition. Now, if $\mathfrak{a}=(f)$ is a principal ideal, then
[…]
    $\Leftrightarrow xf\in I^{F}$
    $\Leftrightarrow x\in I^{F}:(f).$
    Thus, $(I^{F(f)})^{F(f)}=(I^{F}:(f)^{2})$ which is strictly larger than $I^{F(f)}=I^{F}:(f)$ under the given assumptions.
  2. (b)
    Follows immediately from the definition of a non-zerodivisor.
  3. (c)
    Note that as $t>\mu(\mathfrak{a}),$ $\mathfrak{a}^{\lceil tp^{e}\rceil}=\mathfrak{a}^{\lceil(t-1)p^{e}\rceil}\mathfrak{a}^{[p^{e}]}.$ Thus,
    $x\in I^{F\mathfrak{a}^{t}}\Leftrightarrow\forall e\gg 0,x^{p^{e}}\in I^{[p^{e}]}:\mathfrak{a}^{\lceil tp^{e}\rceil}=I^{[p^{e}]}:\mathfrak{a}^{\lceil(t-1)p^{e}\rceil}\mathfrak{a}^{[p^{e}]}\Leftrightarrow x\in(I^{F\mathfrak{a}^{t-1}}:\mathfrak{a})$
    and
    $x\in{}^{\mathfrak{a}^{t}}I^{F}\Leftrightarrow\forall e\gg 0,x^{p^{e}}\in I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil}:\mathfrak{a}^{\lceil tp^{e}\rceil}=(\mathfrak{a}I)^{[p^{e}]}\mathfrak{a}^{\lceil(t-1)p^{e}\rceil}:\mathfrak{a}^{\lceil(t-1)p^{e}\rceil}\mathfrak{a}^{[p^{e}]}\Leftrightarrow x\in(^{\mathfrak{a}^{t-1}}(\mathfrak{a}I)^{F}:\mathfrak{a}).$
  4. (d)
    If $\mathfrak{a}^{n}=0$ for some $n$, then for all $e\gg 0$ and $x\in R$, we have $x^{p^{e}}\cdot 0\subset I^{[p^{e}]}\cdot 0=(0).$
3.4. Comparing the new closures
In this subsection, we compare the different adjusted closures with one another. All statements which involve the adjusted integral closures only are characteristic independent, but whenever the adjusted tight or Frobenius closures are mentioned, we tacitly assume the ambient ring is of prime characteristic $p>0$. Furthermore, in any statement where we invoke the adjusted integral closures, we assume $I\cap R^{\circ}$ and $\mathfrak{a}\cap R^{\circ}$ are nonempty to ensure the adjusted closures are ideals.
By taking $c=1$ in the definition of the (adjusted) $\mathfrak{a}^{t}$-tight closures, we can easily see $I^{F\mathfrak{a}^{t}}\subset I^{*\mathfrak{a}^{t}}$ and $\,{}^{\mathfrak{a}^{t}}I^{F}\subset\,^{\mathfrak{a}^{t}}I^{*}$. Similarly, the adjusted $\mathfrak{a}^{t}$-tight closures are inside the corresponding adjusted $\mathfrak{a}^{t}$-integral closure. In general, we have the following inclusions of ideal closures, fixing $I$, $\mathfrak{a}$, and $t>0$, and assuming $I\cap R^{\circ}$ and $\mathfrak{a}\cap R^{\circ}$ are nonempty.
[…]
Later in Section 4 (Proposition 4.12), we see that the inclusions above are generally strict. We note that each of the adjusted closures are insensitive to replacing the base ideal $I$ with its corresponding unadjusted closure.
Proposition 3.17.
Let $I,\mathfrak{a}\subset R$ be ideals. Then, we have the following equalities.
  1. (a)
    $(I^{F})^{F\mathfrak{a}^{t}}=I^{F\mathfrak{a}^{t}}$ and ($\,{}^{\mathfrak{a}^{t}}(I^{F})^{F}=\,^{\mathfrak{a}^{t}}I^{F}$, and
  2. (b)
    $(I^{*})^{*\mathfrak{a}^{t}}=I^{*\mathfrak{a}^{t}}$ and ($\,{}^{\mathfrak{a}^{t}}(I^{*})^{*}=\,^{\mathfrak{a}^{t}}I^{*}$, and
  3. (c)
    $\overline{J}^{\mathfrak{a}^{t}}=\overline{I}^{\mathfrak{a}^{t}}$, where $J=\overline{I}$.
Proof.
Since (${}^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{F})^{F}={}^{\mathfrak{a}^{t}}I^{F}$ and (${}^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{*})^{*}={}^{\mathfrak{a}^{t}}I^{*},$ it is easy to check the equalities using the inclusions $I\subset I^{F}\subset{}^{\mathfrak{a}^{t}}I^{F}$ and $I\subset I^{*}\subset{}^{\mathfrak{a}^{t}}I^{*},$ so we turn to the Hara-Yoshida versions.
  1. (a)
    We will show $(I^{F})^{F\mathfrak{a}^{t}}=I^{F\mathfrak{a}^{t}}$. As $I\subset I^{F}$, we have $I^{F\mathfrak{a}^{t}}\subset(I^{F})^{F\mathfrak{a}^{t}}$. To see the reverse containment, let $x\in(I^{F})^{F\mathfrak{a}^{t}}$. Then $x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(I^{F})^{\left[p^{e}\right]}=I^{\left[p^{e}\right]}$ for all $e$ large, in particular, for $e$ at least the Frobenius test exponent [^1] There is an $e$ so that $(I^{F})^{\left[p^{e}\right]}\subset I^{\left[p^{e}\right]}$, the least such $e$ is called the Frobenius test exponent of $I$. See [Mad19] for survey of Frobenius test exponents. of $I$. Thus, $x\in I^{F\mathfrak{a}^{t}}$.
  2. (b)
    We will show $(I^{*})^{*\mathfrak{a}^{t}}=I^{*\mathfrak{a}^{t}}$. As $I\subset I^{*}$, we have $I^{*\mathfrak{a}^{t}}\subset(I^{*})^{*\mathfrak{a}^{t}}$. To see the reverse containment, let $x\in(I^{*})^{*\mathfrak{a}^{t}}$. Then there exists $c\in R^{\circ}$ such that $cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(I^{*})^{\left[p^{e}\right]}$ for all $e$ large. There is a $d\in R^{\circ}$ such that $d(I^{*})^{\left[p^{e}\right]}\subset I^{\left[p^{e}\right]}$ for all $e\gg 0$. Thus for $e$ large, $cdx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{\left[p^{e}\right]}$ implying that $x\in I^{*\mathfrak{a}^{t}}$.
  3. (c)
    This proof is nearly identical to the tight closure case, using [SH06, Corollary 6.8.12]. ∎
In [HY03, Proposition 1.3(4)], the authors prove that the Hara-Yoshida $\mathfrak{a}$-tight closure of an ideal $I$ is unaffected by replacing $\mathfrak{a}$ with any ideal with the same integral closure. We prove that the same holds for Vraciu $\mathfrak{a}$-tight closure and $\mathfrak{a}$-integral closure as well.
Proposition 3.18.
Let $I,\mathfrak{a},\mathfrak{b}\subset R$ be ideals and let $\mathfrak{b}\subset\mathfrak{a}$ be a reduction. Then ${}^{\mathfrak{a}^{t}}I^{*}=$ ${}^{\mathfrak{b}^{t}}I^{*}$ and $\overline{I}^{\mathfrak{a}^{t}}=\overline{I}^{\mathfrak{b}^{t}}.$
Proof.
As $\mathfrak{b}\subset\mathfrak{a}$ is a reduction, there exists $k$ such that $\mathfrak{b}^{m}\mathfrak{a}^{k}=\mathfrak{a}^{m+k},$ for all $m\geq 1.$ As $\mathfrak{a}\cap R^{\circ}\neq\varnothing,$ pick $d\in\mathfrak{a}^{k}\cap R^{\circ}.$
We first consider the adjusted tight closures. Let $x\in$ ${}^{\mathfrak{b}^{t}}I^{*}.$ Then there exists $c\in R^{\circ}$ such that $cx^{p^{e}}\mathfrak{b}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\mathfrak{b}^{\lceil tp^{e}\rceil}$ for all $e$ large. Thus, for all $e$ large
$cdx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil+k}\subset cx^{p^{e}}\mathfrak{b}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\mathfrak{b}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil},$
implying that $x\in$ ${}^{\mathfrak{a}^{t}}I^{*}.$ This proves that ${}^{\mathfrak{b}^{t}}I^{*}\subset$ ${}^{\mathfrak{a}^{t}}I^{*}.$
Now, let $x\in$ ${}^{\mathfrak{a}^{t}}I^{*}.$ Then there exists $c\in R^{\circ}$ such that $cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil}$ for all $e$ large. For $e$ large, we have
$cdx^{p^{e}}\mathfrak{b}^{\lceil tp^{e}\rceil}\subset cdx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset dI^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil+k}\subset I^{[p^{e}]}\mathfrak{b}^{\lceil tp^{e}\rceil}$
implying that $x\in$ ${}^{\mathfrak{b}^{t}}I^{*}.$
We now turn to the adjusted integral closures. Since $\mathfrak{b}\subset\mathfrak{a},$ from 3.3(3), we know that $\overline{I}^{\mathfrak{a}^{t}}\subset\overline{I}^{\mathfrak{b}^{t}}.$ Let $x\in\overline{I}^{\mathfrak{b}^{t}}.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{b}^{\lceil tn\rceil}\subset I^{n}$ for all $n$ large. Thus for $n$ large, we have
$cdx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset cx^{n}\mathfrak{a}^{\lceil tn\rceil+k}\subset cx^{n}\mathfrak{b}^{\lceil tn\rceil}\subset I^{n}$
implying that $x\in\overline{I}^{\mathfrak{a}^{t}}.$
One of the most important theorems involving the ordinary tight and integral closures of ideals is the Briançon-Skoda theorem, see [HH90, Theorem 5.4]. A similar version involving the Frobenius closure is proved in [HH94, Discussion 7.6(c)] or [KZ19, Theorem 2.2]. We note that these adjusted closure operations also satisfy a similar theorem.
Theorem 3.19 (A Briançon-Skoda theorem for the adjusted closures).
Let $\mathfrak{a},I$ be ideals in a ring $R$. Let $I$ be generated by $d$ elements and let $t\in\mathbb{R}_{>0}$.
  1. (a)
    Then $\overline{I^{d+m}}^{\mathfrak{a}^{t}}\subset$ $(I^{m+1})^{*\mathfrak{a}^{t}}$ for all $m\geq 0,$ and
  2. (b)
    $\overline{I^{d+m}}^{\mathfrak{a}^{t}}\subset(I^{m})^{F\mathfrak{a}^{t}},$ for all $m\geq 1.$
Proof.
  1. (a)
    If $d=0,$ that is, if $I$ is the zero ideal, then by definition,
    $\overline{(0)^{m}}^{\mathfrak{a}^{t}}=\overline{(0)}^{\mathfrak{a}^{t}}=(0)^{*\mathfrak{a}^{t}}=((0)^{m+1})^{*\mathfrak{a}^{t}}.$
    Now let $I$ be a non-zero ideal and let $x\in\overline{I^{d+m}}^{\mathfrak{a}^{t}}.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n},$ for all $n$ large. Write $n$ as $p^{e}.$ From [Tay21, Lemma 2.1], we know that $I^{(d+m)p^{e}}\subset(I^{[p^{e}]})^{m+1}.$ Thus, we have $cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(I^{m+1})^{[p^{e}]}$ for all $e$ large so that $x\in(I^{m+1})^{*\mathfrak{a}^{t}}$.
  2. (b)
    Let $\sqrt{(0)}$ denote the nilradical of $R.$ By passing to the quotient ring $R/\sqrt{(0)}$, we may assume that $R$ has no embedded primes. In other words, every $c\in R^{\circ}$ is a nonzerodivisor. Let $x\in\overline{I^{d+m}}^{\mathfrak{a}^{t}}.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n},$ for all $n$ large implying that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n}\cap(c)R$. Using Artin-Rees Lemma, there exists $n^{\prime}$ such that for all $n$ large,
    $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n}\cap(c)R\subset cI^{(d+m)n-n^{\prime}}.$
    As $c$ is a nonzerodivisor, we get $x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n-n^{\prime}}$ for all $n$ large. Write $n$ as $p^{e}.$ From [Tay21, Lemma 2.1], we know that $I^{(d+m)p^{e}-n^{\prime}}\subset(I^{[p^{e}]})^{\lceil m-(n^{\prime}/p^{e})\rceil}=(I^{[p^{e}]})^{m},$ for all $e$ large. Thus, $x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(I^{m})^{[p^{e}]}$ for all $e$ large implying that $x\in(I^{m})^{F\mathfrak{a}^{t}}.$
    Generally, if $R$ is not reduced, then from the previous case, we know that for all $m\geq 1,$
    $\overline{I^{d+m}}^{\mathfrak{a}^{t}}=\overline{I^{d+m}}^{\mathfrak{a}^{t}}+\sqrt{(0)}\subset(I^{m})^{F\mathfrak{a}^{t}}+\sqrt{(0)}=(I^{m})^{F\mathfrak{a}^{t}},$
    where the last equality follows from the observation that $\sqrt{(0)}=(0)^{F}\subset(I^{m})^{F}.$
Remark 3.20.
First note that when $\mathfrak{a}=R,$ the above result recovers the usual Briançon-Skoda theorem for tight closure. Further, a similar Briançon-Skoda theorem holds for the Vraciu $\mathfrak{a}^{t}$-integral and tight closures, but since ${}^{\mathfrak{a}^{t}}\overline{I}=\overline{I}$, the result is true simply by the usual Briançon-Skoda; ${}^{\mathfrak{a}^{t}}\overline{I^{m+n}}=\overline{I^{m+n}}\subset(I^{m+1})^{*}\subset$ (${}^{\mathfrak{a}^{t}}(I^{m+1})^{*}.$
To conclude this section, we say a few words about these adjusted closures in the graded setting and prove a graded version of the Briançon-Skoda Theorem.
Setting 3.21.
Let $R=\oplus_{n\geq 0}R_{n}$ be a reduced, finitely generated graded algebra over a field $R_{0}=\mathbb{F}$, and further let $R$ be of positive Krull dimension. Write $R_{+}=\oplus_{n>0}R_{n}$ for the unique maximal homogeneous ideal of $R,$ and let $y_{1},\ldots,y_{s}$ be algebra generators for $R$ over $R_{0}$, with degrees $\beta_{1},\ldots,\beta_{s}$ respectively. Set $\beta=\max(\beta_{i})$ and $\beta^{\prime}=\min(\beta_{i}).$ Finally, we let $\mathfrak{a}$ be a homogeneous $R_{+}$-primary ideal, so that $R^{k}_{+}\subset\mathfrak{a}\subset R^{l}_{+}$ for some integers $l\leq k.$
Proposition 3.22.
Given 3.21, let $I=(f_{1},\ldots,f_{d})$ be a homogeneous ideal, and $x\in R_{\geq N}$ with $N=\beta k-\beta^{\prime}l+\max(\deg(f_{i})).$
  1. (a)
    Then, $x\in\overline{I}^{\mathfrak{a}}$ if and only if $x\in$ ${}^{\mathfrak{a}}\overline{I}$ = $\overline{I}$, and
  2. (b)
    $x\in I^{F\mathfrak{a}}$ if and only if $x\in$ ${}^{\mathfrak{a}}I^{F}.$
Proof.
The proof of the second claim is a quick adaptation of the proof of [Vra08, Proposition 2.9], by simply removing the multiplier $c$ from the equations involved. We now turn to the first claim.
Assume that $x\in\overline{I}^{\mathfrak{a}}$ and $\deg(x)\geq N.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{n}\subset I^{n}$ for all $n$ large. Let $h\in\mathfrak{a}^{n}$ be a homogeneous element. Then $\deg(h)\geq\beta^{\prime}ln.$ Write $cx^{n}h=\sum_{|\mathbf{m}|=n}b_{\mathbf{m}}f_{1}^{m_{1}}\cdots f_{d}^{m_{d}}$ where $b_{i}\in R$ are homogeneous elements, so that
$\deg(b_{\mathbf{m}})\geq\deg(c)+n\deg(x)+\deg(h)-\max(\deg(f_{i}))n\geq\beta kn,$
for each $\mathbf{m}.$ This implies that $b_{\mathbf{m}}\in R_{\geq\beta kn}\subset R_{+}^{kn}\subset\mathfrak{a}^{n},$ hence the result. ∎
The Briançon-Skoda Theorem for graded rings and ideals generated by homogeneous system of parameters was proved in [HH94, Theorem 7.7] (and in [Smi97, Proposition 3.3] for homogeneous $R_{+}$-primary ideals). We prove a similar result for the Hara-Yoshida $\mathfrak{a}$-tight closure. Since these ideals are bigger, as expected, more graded components of the ring are contained in the adjusted closure as compared to the ordinary tight closure.
Proposition 3.23 (Graded Briançon-Skoda Theorem for adjusted closures).
Given 3.21, suppose $\operatorname{char}\mathbb{F}=p>0$ and $R$ is locally equidimensional. Let $I=(f_{1},\ldots,f_{d})$ be such that $f_{1},\ldots,f_{d}$ are homogeneous system of parameters, and let $N=\sum_{i=1}^{d}\deg f_{i}-\beta^{\prime}l.$
Suppose $R$ has a parameter test element [^2] Recall an element $c\in R^{\circ}$ is a parameter test element if, for all $e\geq 0$ and all ideals $\mathfrak{q}$ generated by a system of parameters of $R$, $c(\mathfrak{q}^{*})^{\left[p^{e}\right]}\subset\mathfrak{q}^{\left[p^{e}\right]}$., e.g. $\mathbb{F}$ is $F$-finite. Let $x\in R_{\geq N}.$ Then $x\in I^{*\mathfrak{a}}.$ In other words, $R_{\geq N}+I^{F\mathfrak{a}}\subset I^{*\mathfrak{a}}.$
Proof.
We first prove the case in which $\deg f_{i}=\deg f_{1}=\alpha$ for all $1\leq i\leq d.$ Pick $v\in\mathfrak{a}$ a homogeneous element. As $\deg v\geq\beta^{\prime}l,$ we get $\deg(xv)\geq d\alpha.$ Since $R$ is module-finite over $S=\mathbb{F}[f_{1},\ldots,f_{d}],$ $xv$ satisfies an integral equation $(xv)^{m}+b_{1}(xv)^{m-1}+\cdots+b_{m}=0$ where $b_{i}\in S$ for all $i.$ We may assume that $b_{i}$ are homogeneous elements. So $\deg b_{i}=i\cdot\deg(xv)$ implying that
$b_{i}\in R_{i\cdot\deg(xv)}\cap S\subset(f_{1},\ldots,f_{d})^{i\cdot\deg(xv)/\alpha}\subset(f_{1},\ldots,f_{d})^{id}.$
This implies that $xv\in\overline{I^{d}}.$ By definition, there exists $c_{0}\in R^{\circ}$ and $n_{0}$ such that $c_{0}x^{n}v^{n}\in I^{nd},$ for all $n\geq n_{0}.$ Note that $I^{n_{0}}\cap R^{\circ}\neq\varnothing.$ Choose $c_{1}\in I^{n_{0}}\cap R^{\circ}$ and set $c=c_{0}c_{1}.$ Then for all $n\geq 0,$
(3.1) $cx^{n}v^{n}\in I^{nd}.$
We claim that $x\in\overline{(I^{d})}^{\mathfrak{a}}.$ Let $\mathfrak{a}=(v_{1},\ldots,v_{\mu}).$ For $1\leq i\leq\mu,$ from (3.1) we know that there exists $c_{i}\in R^{\circ}$ such that $c_{i}x^{n}v_{i}^{n}\in I^{nd}$ for all $n.$ Then for $n_{1}+\cdots+n_{\mu}=n,$
$\left(\prod_{i=1}^{\mu}c_{i}\right)x^{n}v_{1}^{n_{1}}\cdots v_{\mu}^{n_{\mu}}=\prod_{i=1}^{\mu}\big{(}c_{i}x^{n_{i}}v_{i}^{n_{i}}\big{)}\subset I^{nd},$
implying that $x\in\overline{(I^{d})}^{\mathfrak{a}}\subset I^{*\mathfrak{a}}$ using Theorem 3.19.
Now suppose that not all $f_{i}$’s have the same degree. Choose nonnegative integers $\alpha_{i}$ such that $F_{1}=f_{1}^{\alpha_{1}+1},\ldots,F_{d}=f_{d}^{\alpha_{d}+1}$ have equal degrees, say $D$. Let $x\in R_{N}$ where $N\geq\sum_{i=1}^{d}\deg f_{i}-\beta^{\prime}l.$ Then $xf_{1}^{\alpha_{1}}\cdots f_{d}^{\alpha_{d}}$ has degree at least $dD-\beta^{\prime}l$ and hence by the first case is in $(F_{1},\ldots,F_{d})^{*\mathfrak{a}}.$ By definition, there exists $c\in R^{\circ}$ such that for all $e$ large,
$c(xf_{1}^{\alpha_{1}}\cdots f_{d}^{\alpha_{d}})^{p^{e}}\mathfrak{a}^{p^{e}}\subset(f_{1}^{(\alpha_{1}+1)p^{e}},\ldots,f_{d}^{(\alpha_{d}+1)p^{e}}).$
This implies that
$cx^{p^{e}}\mathfrak{a}^{p^{e}}\subset(f_{1}^{(\alpha_{1}+1)p^{e}},\ldots,f_{d}^{(\alpha_{d}+1)p^{e}}):(f_{1}^{\alpha_{1}}\cdots f_{d}^{\alpha_{d}})^{p^{e}}\subset(f_{1}^{p^{e}},\ldots,f_{d}^{p^{e}})^{*}$
where the last inclusion follows from the colon capturing property (see [Hun98, Theorem 2.3(b)]). Multiplying by a parameter test element $c^{\prime}$, we get $cc^{\prime}x^{p^{e}}\mathfrak{a}^{p^{e}}\subset(f_{1},\ldots,f_{d})^{[p^{e}]}$ for all $e$ large implying that $x\in I^{*\mathfrak{a}}$. ∎
4. Adjusted Ideal Closures in Semigroup Rings
We now consider the class of rings $k[S]$, where $k$ is a field and $S$ is a submonoid of the free abelian group $M=\mathbb{Z}^{d}$ of rank $d$. We have that $M$ naturally sits inside $V=M\otimes_{\mathbb{Z}}\mathbb{R}\cong\mathbb{R}^{d}$ and we can utilize convex geometry in $\mathbb{R}^{d}$ to analyze monomial ideals in $k[S]$. We will briefly describe the tools we need; we refer the reader to books such as [Ful93] or [CLS11] for further background on affine semigroup rings.
Setting 4.1.
Fix a field $k$. Let $M$ be a rank $d$ free abelian group, $S$ a finitely generated submonoid of $M$ such that $S$ generates $M$ as a group, and $V=M\otimes_{\mathbb{Z}}\mathbb{R}$. Let $\sigma^{\vee}=\mathrm{cone}(S)\subset V$, and $\sigma=\mathrm{cone}(\mathbf{c}_{1},\ldots,\mathbf{c}_{n})\subset V^{*}$ the dual cone to $\sigma^{\vee}$. That is, $\mathbf{v}\in\sigma^{\vee}$ if and only if $\mathbf{v}\in V$ and $\langle\mathbf{c}_{i},\mathbf{v}\rangle\geq 0$ for all $i=1,\ldots,n$. Let $\bar{S}=\sigma^{\vee}\cap M$. Thus, $\bar{S}$ is the saturation of $S$ inside $M$, and $\bar{R}=k[\bar{S}]$ is the integral closure of $R=k[S]$ inside $k[M]$. Note that for any ideal $J$ of $R$, $(\operatorname{Exp}(J)+\sigma^{\vee})\cap M=\operatorname{Exp}(J\bar{R})$.
Let $I$ and $\mathfrak{a}$ be monomial ideals of $R$, so that there exist finite sets $\mathcal{A},\mathcal{B}\subset S$ such that $\mathfrak{a}=(x^{\mathbf{a}}\mid\mathbf{a}\in\mathcal{A})$ and $I=(x^{\mathbf{b}}\mid\mathbf{b}\in\mathcal{B})$. Let $A=(\mathbf{a}_{1}|\cdots|\mathbf{a}_{r})$, and let $\Delta=\{\mathbf{u}\in\mathbb{R}_{\geq 0}^{r}\mid|\mathbf{u}|_{1}=1\}$. Let $H=\mathrm{Hull}(\mathbf{a}_{1},\ldots,\mathbf{a}_{r})$ be the convex hull of the columns of $A$, and notice that $H=\{A\mathbf{u}\mid\mathbf{u}\in\Delta\}$.
Lemma 4.2.
Given 4.1, for each $n\in\mathbb{Z}_{\geq 0}$ and $s\in\mathbb{R}_{>0}$, we have
  1. (a)
    $\operatorname{Exp}(\mathfrak{a}^{n})\subset nH+\sigma^{\vee}$, and
  2. (b)
    $sH\subset\operatorname{Exp}(\mathfrak{a}^{\lfloor s-d\rfloor})+\sigma^{\vee}$.
Proof.
  1. (a)
    For any $n\in\mathbb{N}$, $\mathfrak{a}^{n}$ is generated by monomials of the form $A\mathbf{w}$, where $\mathbf{w}\in\mathbb{N}^{r}$ and $|\mathbf{w}|_{1}=n$. Letting $\mathbf{u}=\frac{1}{n}\mathbf{w}$, we have that $\mathbf{u}\in\Delta$ and so $A\mathbf{w}=nA\mathbf{u}\in nH$. Therefore $\operatorname{Exp}(\mathfrak{a}^{n})\subset nH+\sigma^{\vee}.$
  2. (b)
    On the other hand, if $\mathbf{v}\in sH$ and $s>0$, then there exists $\mathbf{u}\in\Delta$ such that $\mathbf{v}=sA\mathbf{u}$. Furthermore, by Carathéodory’s theorem, we may assume that at most $d+1$ of the components of $u$ are nonzero. Let $\mathbf{w}\in\mathbb{N}^{r}$ be given by $w_{i}=\lfloor su_{i}\rfloor$. Now,
    $\mathbf{v}=\mathbf{v}-A\mathbf{w}+A\mathbf{w}=A(s\mathbf{u}-\mathbf{w})+A\mathbf{w}\in A\mathbf{w}+\sigma^{\vee}.$
    Since $|\mathbf{w}|_{1}=\sum_{i}\lfloor su_{i}\rfloor>\sum_{i:u_{i}\neq 0}(su_{i}-1)\geq s-d-1$, we have that $|\mathbf{w}|_{1}\geq\lfloor s-d\rfloor$ and so $A\mathbf{w}\in\operatorname{Exp}(\mathfrak{a}^{\lfloor s-d\rfloor})$. Thus, we have that $sH\subset\operatorname{Exp}(\mathfrak{a}^{\lfloor s-d\rfloor})+\sigma^{\vee}$, and so $sH\cap\bar{S}\subset\operatorname{Exp}(\mathfrak{a}^{\lfloor s-d\rfloor}\bar{R})$.
4.1. Joint Hilbert-Kunz multiplicity
In [Vra08], the author defined the joint Hilbert-Kunz multiplicity, $e_{HK}(\mathfrak{a}^{t};I),$ of two finite colength ideals $I,\mathfrak{a}$ in a Noetherian local ring $(R,\mathfrak{m})$ of dimension $d\geq 1.$ Letting $t>0$ be a fixed real number,
$e_{HK}(\mathfrak{a}^{t};I)=\lim\limits_{q\rightarrow\infty}\frac{1}{q^{d}}\ell\left(\frac{R}{\mathfrak{a}^{\lceil tq\rceil}I^{[q]}}\right).$
It is known that the Hilbert-Kunz multiplicity of finite length ideals in a semigroup ring can be computed as the volume of a finite union of rational polytopes. The following result proves that the same holds for the joint Hilbert-Kunz multiplicity.
Theorem 4.3.
Given 4.1, suppose $k$ is of prime characteristic $p>0$, and let $I$ and $\mathfrak{a}$ be ideals in $R$ of finite colength, and let $t>0$ be a fixed real number. If we set
$\mathcal{P}_{t}=\sigma^{\vee}\setminus(\mathcal{B}+tH+\sigma^{\vee}),$
then $e_{HK}(\mathfrak{a}^{t};I)=\operatorname{vol}\ \overline{\mathcal{P}_{t}}$, where $\overline{\mathcal{P}_{t}}$ is the closure of $\mathcal{P}_{t}$ in the Euclidean topology on $\mathbb{R}^{d}$ and $\operatorname{vol}$ denotes the usual Euclidean volume on $\mathbb{R}^{d}$.
Proof.
First note that it is enough to prove the statement when $R$ is a normal ring, as if $R$ is not normal, then from [Vra08, Lemma 3.3] we get that $e_{HK}(\mathfrak{a}^{t};I)=e_{HK}(\mathfrak{a}^{t}\bar{R};I\bar{R})$ as $\overline{R}$ is a module-finite extension of $R$.
Let $g_{1},\ldots,g_{n}$ be the generators of $\sigma^{\vee}$ and let $F_{1},\ldots,F_{\alpha}$ be the cones divided by all the hyperplanes spanned by $g_{1},\ldots,g_{n}$ and $\mathbf{b}+t\mathbf{a}_{i}$, where $b\in\mathcal{B}$ and $1\leq i\leq r.$ Then for each $\mathbf{b}\in\mathcal{B},$ the closure of the complement of $(\mathbf{b}+tH+\sigma^{\vee})$ in $\sigma^{\vee}$ is a finite union of convex sets $\overline{F_{i}\cap(\mathbf{b}+tH+\sigma^{\vee})^{c}}.$ Hence,
$\mathcal{P}_{t}=\sigma^{\vee}\bigcap\left(\bigcap_{\mathbf{b}\in\mathcal{B}}\overline{\big{(}\mathbf{b}+tH+\sigma^{\vee}\big{)}^{c}}\right)=\bigcap_{\mathbf{b}\in\mathcal{B}}\left(\bigcup_{i=1}^{\alpha}\overline{F_{i}\cap\big{(}\mathbf{b}+tH+\sigma^{\vee}\big{)}^{c}}\right)$
is a finite union of rational polytopes, say $P_{t,1},\ldots,P_{t,\beta},$ containing the origin such that for all $i\neq j,$ $P_{t,i}\cap P_{t,j}$ is a rational polytope of dimension $<d.$ Using the same arguments as in the proof of [Eto02, Theorem 2.2], we get
(4.1) $\operatorname{vol}\overline{\mathcal{P}_{t}}=\sum_{i=1}^{\beta}\operatorname{vol}P_{t,i}=\sum_{i=1}^{\beta}\lim\limits_{q\rightarrow\infty}\frac{\mid qP_{t,i}\cap\mathbb{Z}^{d}\mid}{q^{d}}=\lim\limits_{q\rightarrow\infty}\frac{\mid q\mathcal{P}_{t}\cap\mathbb{Z}^{d}\mid}{q^{d}}.$
In order to compute the joint Hilbert-Kunz multiplicity, we observe that if $x^{\mathbf{v}}\in\mathfrak{a}^{\lceil tq\rceil}I^{[q]}$, then $\mathbf{v}\in A\mathbf{w}+q\mathcal{B}+\sigma^{\vee}$, where $\mathbf{w}\in\mathbb{N}^{r}$ and $|\mathbf{w}|_{1}=\lceil tq\rceil$, so we have
$\mathbf{v}\in\lceil tq\rceil H+q\mathcal{B}+\sigma^{\vee}\subset q(tH+\mathcal{B}+\sigma^{\vee})=\sigma^{\vee}\setminus q\mathcal{P}_{t}.$
Therefore, if $\mathbf{v}\in q\mathcal{P}_{t}$, then $x^{\mathbf{v}}\notin\mathfrak{a}^{\lceil tq\rceil}I^{[q]}$, and so $\ell(R/\mathfrak{a}^{\lceil tq\rceil}I^{[q]})\geq|q\mathcal{P}_{t}\cap\mathbb{Z}^{d}|$.
On the other hand, suppose that $s>t$, and $q$ is large enough that $\lfloor sq-d\rfloor\geq\lceil tq\rceil$. If $\mathbf{v}\in S\setminus q\mathcal{P}_{s}$, then
$\mathbf{v}\in q\mathcal{B}+sqH+\sigma^{\vee}\subset$ $\operatorname{Exp}(I^{[q]})+\operatorname{Exp}(\mathfrak{a}^{\lfloor sq-d\rfloor})+\sigma^{\vee}$
$=$ $\operatorname{Exp}(\mathfrak{a}^{\lfloor sq-d\rfloor}I^{[q]})+\sigma^{\vee}\subset\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil}I^{[q]})+\sigma^{\vee}.$
Therefore $x^{\mathbf{v}}\in\mathfrak{a}^{\lceil tq\rceil}I^{[q]}$. Hence $\ell(R/\mathfrak{a}^{\lceil tq\rceil}I^{[q]})\leq|q\mathcal{P}_{s}\cap\mathbb{Z}^{d}|$.
Thus, we have that, for any $s>t$ and $q\gg 0$,
$|q\mathcal{P}_{t}\cap\mathbb{Z}^{d}|\leq\ell(R/\mathfrak{a}^{\lceil tq\rceil}I^{[q]})\leq|q\mathcal{P}_{s}\cap\mathbb{Z}^{d}|$
and hence, for any $s>t$,
$\lim_{q\to\infty}\frac{|q\mathcal{P}_{t}\cap\mathbb{Z}^{d}|}{q^{d}}\leq e_{HK}(\mathfrak{a}^{t};I)\leq\lim_{q\to\infty}\frac{|q\mathcal{P}_{s}\cap\mathbb{Z}^{d}|}{q^{d}}.$
From (4.1), we get $\operatorname{vol}\overline{\mathcal{P}_{t}}\leq e_{HK}(\mathfrak{a}^{t};I)\leq\lim_{s\to t^{+}}\operatorname{vol}\overline{\mathcal{P}_{s}}=\operatorname{vol}\overline{\mathcal{P}_{t}}.$
4.2. The adjusted closures in affine semigroup rings
In this subsection we show that the calculation of the adjusted ideal closures in affine semigroup rings are computable in terms of the convex geometry of the associated exponent sets.
Theorem 4.4.
Given 4.1, suppose $k$ is of prime characteristic $p>0$, and let $t>0$ and $x^{\mathbf{v}}\in R$ be a monomial.
  1. (a)
    If $x^{\mathbf{v}}\in I^{*\mathfrak{a}^{t}}$, in particular if $x^{\mathbf{v}}\in I^{F\mathfrak{a}^{t}}$, then $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$.
  2. (b)
    If $x^{\mathbf{v}}\in{{}^{\mathfrak{a}^{t}}I^{*}}$, in particular if $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}I^{F}$ then $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$.
Proof.
Suppose that there exists $x^{\mathbf{c}}\in R$, $s\geq 0$ such that for $q\gg 0$, $x^{\mathbf{c}}x^{q\mathbf{v}}\mathfrak{a}^{\lceil tq\rceil}\subset I^{[q]}\mathfrak{a}^{\lceil sq\rceil}$. That is,
$\mathbf{c}+q\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil})\subset q\mathcal{B}+\operatorname{Exp}(\mathfrak{a}^{\lceil sq\rceil})+\sigma^{\vee}.$
This implies that
$\mathbf{c}+q\mathbf{v}+\lceil tq+d\rceil H$ $\subset\mathbf{c}+q\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil})+\sigma^{\vee}$
$\subset q\mathcal{B}+\operatorname{Exp}(\mathfrak{a}^{\lceil sq\rceil})+\sigma^{\vee}$
$\subset q\mathcal{B}+\lceil sq\rceil H+\sigma^{\vee}.$
Since $H\subset\sigma^{\vee}$, $\lceil sq\rceil H+\sigma^{\vee}\subset sqH+\sigma^{\vee}$. Therefore, dividing through by $q$, we have that for $q\gg 0$,
$\frac{1}{q}\mathbf{c}+\mathbf{v}+\frac{\lceil tq+d\rceil}{q}H\subset\mathcal{B}+sH+\sigma^{\vee}.$
Now, since $\mathcal{B}+sH+\sigma^{\vee}$ is a closed set, for any $\mathbf{h}\in H$ and $q\geq 1$,
$\mathbf{v}+t\mathbf{h}=\lim_{q\to\infty}\left(\frac{1}{q}\mathbf{c}+\mathbf{v}+\frac{\lceil tq+d\rceil}{q}\mathbf{h}\right)\in\mathcal{B}+sH+\sigma^{\vee}.$
Therefore $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$.
Taking $s=0$ shows that if $x^{\mathbf{v}}\in I^{*\mathfrak{a}^{t}}$, then $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$, and the same holds if $x^{\mathbf{v}}\in I^{F\mathfrak{a}^{t}}$, since $I^{F\mathfrak{a}^{t}}\subset I^{*\mathfrak{a}^{t}}$. Taking $s=t$ shows that if $x^{\mathbf{v}}\in{{}^{\mathfrak{a}^{t}}I^{*}}$, then $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$, and the same holds if $x\in{}^{\mathfrak{a}^{t}}I^{F}$, since ${}^{\mathfrak{a}^{t}}I^{F}\subset{{}^{\mathfrak{a}^{t}}I^{*}}$. ∎
Theorem 4.5.
Given 4.1, suppose $k$ is of prime characteristic $p>0$, and let $t>0$ and $x^{\mathbf{v}}\in R$ be a monomial.
  1. (a)
    If $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$, then $x^{\mathbf{v}}\in(I\bar{R})^{F\mathfrak{a}^{t}}$ and $x^{\mathbf{v}}\in(I\bar{R})^{*\mathfrak{a}^{t}}$.
  2. (b)
    If $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$, then $x^{\mathbf{v}}\in{{}^{\mathfrak{a}^{t}}(I\bar{R})^{*}}$.
  3. (c)
    If $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ for some $s>t$, then $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}(I\bar{R})^{F}$.
Proof.
Let $s\geq 0$ be a real number, and suppose that $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$. Now, for any $q\geq 1$, we have that
$q\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil})\subset q\mathbf{v}+tqH+\sigma^{\vee}\subset q\mathcal{B}+sqH+\sigma^{\vee}\subset q\mathcal{B}+\operatorname{Exp}(\mathfrak{a}^{\lfloor sq-d\rfloor})+\sigma^{\vee}.$
Note that $(q\mathcal{B}+\operatorname{Exp}(\mathfrak{a}^{\lfloor sq-d\rfloor})+\sigma^{\vee})\cap\bar{S}=\operatorname{Exp}(I^{[q]}\mathfrak{a}^{\lfloor sq-d\rfloor}\bar{R})$.
Suppose $s=0$, so that we have $q\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil})\subset\operatorname{Exp}(I^{[q]}\bar{R})$, and so $x^{q\mathbf{v}}\mathfrak{a}^{\lceil tq\rceil}\subset I^{[q]}\bar{R}$. Hence, if $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$, then $x^{\mathbf{v}}\in(I\bar{R})^{F\mathfrak{a}^{t}}\subset(I\bar{R})^{*\mathfrak{a}^{t}}$.
Suppose $s=t$, and choose $x^{\mathbf{c}}\in\mathfrak{a}^{d+1}$. Now, for $q\geq 1$, we have that
$\mathbf{c}+q\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil})\subset\operatorname{Exp}(x^{\mathbf{c}}I^{[q]}\mathfrak{a}^{\lfloor tq-d\rfloor})\subset\operatorname{Exp}(I^{[q]}\mathfrak{a}^{\lfloor tq+1\rfloor})\subset\operatorname{Exp}(I^{[q]}\mathfrak{a}^{\lceil tq\rceil}).$
Therefore $x^{\mathbf{c}}x^{q\mathbf{v}}\mathfrak{a}^{\lceil tq\rceil}\subset I^{[q]}\mathfrak{a}^{\lceil tq\rceil}$ for all $q\geq 1$. Hence, if $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$, then $x^{\mathbf{v}}\in{{}^{\mathfrak{a}^{t}}(I\bar{R})^{*}}$.
Suppose $s>t$, so that for $q\gg 0$, we have that $sq-d\geq\lceil tq\rceil$, and so $q\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil})\subset\operatorname{Exp}(I^{[q]}\mathfrak{a}^{\lfloor sq-d\rfloor}\bar{R})\subset\operatorname{Exp}(I^{[q]}\mathfrak{a}^{\lceil tq\rceil}\bar{R})$. Hence, if $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ for some $s>t$, then $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}(I\bar{R})^{F}$. ∎
Corollary 4.6.
Given 4.1, if $k$ is of prime characteristic $p>0$ and $R=k[S]$ is a normal semigroup ring, then for a monomial $x^{\mathbf{v}}\in R$, we have
  1. (a)
    $x^{\mathbf{v}}\in I^{F\mathfrak{a}^{t}}$ if and only if $x^{\mathbf{v}}\in I^{*\mathfrak{a}^{t}}$ if and only if $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$,
  2. (b)
    $x^{\mathbf{v}}\in{{}^{\mathfrak{a}^{t}}I^{*}}$ if and only if $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$,
  3. (c)
    if $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}I^{F}$ then $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$, and
  4. (d)
    if there exists $s>t$ such that $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$, then $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}I^{F}$.
Remark 4.7.
The converse of Corollary 4.6(c) does not hold in general, that is, the $\mathfrak{a}^{t}$-tight and Frobenius closures may differ. For example, let $k$ be a field of odd prime characteristic $p$, $R=k[x,y]$, $I=\mathfrak{a}=(x^{2},y^{2})$, and $t=1$, so $tH=\mathrm{Hull}((2,0),(0,2))$. Then, $(1,1)+tH=\mathrm{Hull}((3,1),(1,3))$, and $\{(2,0),(0,2)\}+tH=\mathrm{Hull}((4,0),(0,4))$. Therefore $(1,1)+tH\subset\{(2,0),(0,2)\}+tH$, demonstrating an $\mathfrak{a}^{t}$-tight closure relation. However, $xy\notin{}^{\mathfrak{a}^{t}}I^{F}$; for any power $q$ of $p$, we have that $x^{3q}y^{q}\in(xy)^{q}\mathfrak{a}^{q}$. But, this monomial is not in $I^{[q]}\mathfrak{a}^{q}$, since all generators of the latter are of degree $4q$ but have only even exponents.
Also, the normality condition in Corollary 4.6 is necessary for the implications coming from Theorem 4.5. To see this, let $R=k[x^{2},x^{3},y,xy]$. Note that $R$ contains all monomials in $k[x,y]$ except $x$. Now, let $I=(x^{2}y)$, $\mathfrak{a}=(x^{2},x^{3})$, and $t=1$. This makes $H=\mathrm{Hull}((2,0),(3,0))$, and so
$(0,1)+tH=\mathrm{Hull}((2,1),(3,1))\subset(2,1)+\sigma^{\vee}=\operatorname{Exp}(I)+\sigma^{\vee}.$
Therefore, if Corollary 4.6 were to hold without the normality assumption, we would have that $y\in I^{F\mathfrak{a}^{t}}$. However, for any power $q$ of $p$ we have that
$y^{q}\mathfrak{a}^{q}=(x^{2a+3b}y^{q}\mid a+b=q)=(x^{2q}y^{q},x^{2q+1}y^{q})$
but $I^{[q]}=(x^{2q}y^{q})$. Since $x\notin R$, $x^{2q+1}y^{q}\notin I^{[q]}$. Therefore $y\notin I^{F\mathfrak{a}^{t}}$.
Independent of the characteristic of $k$, we also get similar convex geometry statements for the $\mathfrak{a}^{t}$-integral closure.
Theorem 4.8.
Given 4.1, if $R=k[S]$ is a normal semigroup ring and $x^{\mathbf{v}}\in R$ is a monomial, then $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$ if and only if $\mathbf{v}+tH\subset\operatorname{Hull}(\operatorname{Exp}I)$.
Proof.
The proof is similar to Theorem 4.5, but we provide it for completeness. For simplicity, write $E=\operatorname{Hull}(\operatorname{Exp}I)$.
First, suppose $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$. So, there is a $\mathbf{c}\in S$ with $x^{\mathbf{c}}x^{n\mathbf{v}}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for all $n\gg 0$. In vector form, this is equivalent to $\mathbf{c}+n\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil nt\rceil})\subset\operatorname{Exp}(I^{n})$. By Lemma 4.2, we have
$\mathbf{c}+n\mathbf{v}+\lceil tn+d\rceil H\subset$ $\mathbf{c}+n\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tn\rceil})+\sigma^{\vee}$
$\subset$ $\operatorname{Exp}(I^{n})+\sigma^{\vee}$
$\subset$ $nE+\sigma^{\vee}=nE$
since $nE+\sigma^{\vee}=nE$. Then, given $\mathbf{c}+n\mathbf{v}+\lceil tn+d\rceil H\subset nE$, we divide both sides by $n$, viewed as subsets of $V=M\otimes_{\mathbb{Z}}\mathbb{R}$ to get
$\dfrac{\mathbf{c}}{n}+\mathbf{v}+\dfrac{\lceil tn+d\rceil}{n}H\subset E.$
As $E$ is a closed subset of $\mathbb{R}^{d}$ in the usual Euclidean topology, we may take the limit in $n$ and stay inside $E$, that is, $\mathbf{v}+tH\subset E$. Notably, this direction of the theorem does not require that $R$ be normal.
Now, suppose $\mathbf{v}+tH\subset E$, and we want to show $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$. For some $n\in\mathbb{N}$, let $\mathbf{a}\in\operatorname{Exp}(\mathfrak{a}^{\lceil nt\rceil})$ be such that $x^{\mathbf{a}}$ is a nonzero monomial generator of $\mathfrak{a}^{\lceil tn\rceil}$. Then, there is a $\mathbf{w}$ in $\mathbb{N}^{r}$ such that $|\mathbf{w}|_{1}=\lceil tn\rceil$ and $\mathbf{a}=A\mathbf{w}$. Also, fix $\mathbf{c}\in\operatorname{Exp}(I^{\ell})$, for $\ell=\mu(I)-1$. We then have
$\mathbf{c}+n\mathbf{v}+\mathbf{a}=$ $n\left(\dfrac{\mathbf{c}}{n}+\mathbf{v}+\dfrac{\mathbf{a}}{n}\right)$
$=$ $n\left(\dfrac{\mathbf{c}}{n}+\mathbf{v}+\dfrac{1}{n}A\mathbf{w}\right)$
$=$ $n\left(\dfrac{\mathbf{c}}{n}+\mathbf{v}+\dfrac{t}{\lceil tn\rceil}A\mathbf{w}+\left(\dfrac{1}{n}-\dfrac{t}{\lceil tn\rceil}\right)A\mathbf{w}\right).$
Notably, the term $\mathbf{v}+(t/\lceil tn\rceil)A\mathbf{w}$ is inside $\mathbf{v}+tH\subset E$ and the term $(1/n-t/\lceil tn\rceil)A\mathbf{w}$ is in $\sigma^{\vee}$ since $t/\lceil tn\rceil<t/tn=1/n$. Thus, $\mathbf{c}+n\mathbf{v}+\mathbf{a}$ is inside $\mathbf{c}+nE+\sigma^{\vee}\subset\mathbf{c}+nE\subset\mathbf{c}+\operatorname{Exp}(\overline{I^{n}})$. Finally, by the Briançon-Skoda theorem, $\overline{I^{n}}\subset I^{n-\ell}$ for all $n$. This shows $\mathbf{c}+\operatorname{Exp}(I^{n-\ell})\subset\operatorname{Exp}(I^{n})$ by our choice of $\mathbf{c}$, demonstrating that $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$, as required. ∎
Remark 4.9.
We note that the $\mathfrak{a}^{t}$-tight and integral closure calculations above are entirely independent of characteristic. However, below we demonstrate characteristic dependence in the Frobenius closures.
4.3. Computed Examples
Given 4.1, let $C=(\mathbf{c}_{1}|\mathbf{c}_{2}|\cdots|\mathbf{c}_{n})$ be the matrix whose columns are the vectors $\mathbf{c}_{i}$, which means that a vector $\mathbf{v}\in V$ is in $\sigma^{\vee}$ if and only if $C^{T}\mathbf{v}\geq\mathbf{0}$ componentwise.
For any $s\geq 0$, $t>0$, and $\mathbf{v}\in S$, $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ if and only if for all $\mathbf{u}\in\Delta$ there exists $\mathbf{w}\in\Delta$ and $i$ such that $\mathbf{v}+tA\mathbf{u}\subset\mathbf{b}_{i}+sA\mathbf{w}+\sigma^{\vee}$, that is $C^{T}(\mathbf{v}-\mathbf{b}_{i}+A(t\mathbf{u}-s\mathbf{w}))\geq\mathbf{0}.$
4.3.1. The case $\mu(\mathfrak{a})=2$
Suppose that $A=(\mathbf{a}_{1}\mid\mathbf{a}_{2})$, and hence $\Delta=\{(1-\gamma,\gamma)^{T}\mid\gamma\in[0,1]\}$. Now, for $\mathbf{u}=(1-\gamma,\gamma)^{T}\in\Delta$, $\mathbf{w}=(1-\delta,\delta)^{T}\in\Delta$, and $t,s\geq 0$, we have that
$A(t\mathbf{u}-s\mathbf{w})=(\mathbf{a}_{1}\mid\mathbf{a}_{2})\left(\begin{array}[]{c}t-s-(t\gamma-s\delta)\\ t\gamma-s\delta\end{array}\right)=(t-s)\mathbf{a}_{1}+(t\gamma-s\delta)(\mathbf{a}_{2}-\mathbf{a}_{1}).$
Now,
$C^{T}(\mathbf{v}-\mathbf{b}_{i}+A(t\mathbf{u}-s\mathbf{w}))\geq\mathbf{0}$
$\iff$ $C^{T}A(t\mathbf{u}-s\mathbf{w}))\geq C^{T}\mathbf{b}_{i}-C^{t}\mathbf{v}$
$\iff$ $\forall j,\;\langle\mathbf{c}_{j},(t-s)\mathbf{a}_{1}+(t\gamma-s\delta)(\mathbf{a}_{2}-\mathbf{a}_{1})\rangle\geq\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle$
$\iff$ $\forall j,\;(t\gamma-s\delta)\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle\geq\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle.$
Now for all $j$ with $\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle>0$, the condition above becomes
(4.2) $t\gamma-s\delta\geq\frac{\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle}{\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle}.$
For each $i$, let
$m_{i}(\mathbf{v})=\max\left\{\frac{\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle}{\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle}\mid j:\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle>0\right\}.$
For all $j$ with $\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle<0$, the condition above becomes
(4.3) $t\gamma-s\delta\leq\frac{\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle}{\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle}.$
For each $i$, let
$M_{i}(\mathbf{v})=\min\left\{\frac{\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle}{\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle}\mid j:\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle<0\right\}.$
For any $j$ with $\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle=0$, the condition is
(4.4) $\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle\leq 0\quad\iff\quad\langle\mathbf{c}_{j},\mathbf{v}\rangle\geq\langle\mathbf{c}_{j},\mathbf{b}_{i}-(t-s)\mathbf{a}_{1}\rangle.$
This inequaltiy does not depend on $\gamma$ or $\delta$, and so is a necessary condition on any $\mathbf{v}$ with $\mathbf{v}+tA\mathbf{u}\in\mathcal{B}+sH+\sigma^{\vee}$. For the remainder of this section we assume that this inequality holds on the vectors $\mathbf{v}$ we consider.
Thus, we have that for a particular $\mathbf{u}=(1-\gamma,\gamma)$, that $\mathbf{v}+tA\mathbf{u}\in\mathcal{B}+sH+\sigma^{\vee}$ if and only if there exists $i$ and $\delta\in[0,1]$ such that $m_{i}(\mathbf{v})\leq t\gamma-s\delta\leq M_{i}(\mathbf{v})$. So long as $m_{i}(\mathbf{v})\leq M_{i}(\mathbf{v})$, it will be possible to find a $\delta$ making this inequality true if and only if $m_{i}(\mathbf{v})\leq t\gamma\leq M_{i}(\mathbf{v})+s$. Thus, we have proved the following.
Theorem 4.10.
Given 4.1, let $C=(\mathbf{c}_{1}|\mathbf{c}_{2}|\cdots|\mathbf{c}_{n})$, and assume that $\mathfrak{a}$ is 2-generated, i.e. $A=(\mathbf{a}_{1}\mid\mathbf{a}_{2})$. Let $t>0$ and $s\geq 0$. For each $i$, let $m_{i}:S\to\mathbb{R}$ and $M_{i}:S\to\mathbb{R}$ be defined by
$m_{i}(\mathbf{v})$ $=\max\left\{\frac{\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle}{\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle}\mid j:\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle>0\right\},$
$M_{i}(\mathbf{v})$ $=\min\left\{\frac{\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle}{\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle}\mid j:\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle<0\right\}.$
Now, for any $\mathbf{v}\in S$, we have that $\mathbf{v}+tA\mathbf{u}\in\mathcal{B}+sH+\sigma^{\vee}$ if and only if, for all $j$ with $\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle=0$ we have that $\langle\mathbf{c}_{j},\mathbf{v}\rangle\geq\langle\mathbf{c}_{j},\mathbf{b}_{i}-(t-s)\mathbf{a}_{1}\rangle$, and
$[0,1]\subset\bigcup_{i:m_{i}(\mathbf{v})\leq M_{i}(\mathbf{v})}\left[\frac{m_{i}(\mathbf{v})}{t},\frac{M_{i}(\mathbf{v})+s}{t}\right].$
In the example below, the statements about the adjusted integral closures are independent of the characteristic of $k$, but the statements involving the joint Hilbert-Kunz multiplicity and the adjusted tight and Frobenius closures requires that $k$ is of prime characteristic $p>0$.
Example 4.11.
Let $R=k[x,xy,x^{2}y^{3}]=k[S]$, where $S=\sigma^{\vee}\cap\mathbb{Z}^{2}$ (notably, a normal semigroup) and $\sigma^{\vee}$ is the cone generated by $(1,0)$ and $(2,3)$, which is dual to the cone $\sigma$ generated by $\mathbf{c}_{1}=(0,1)$ and $\mathbf{c}_{2}=(3,-2)$. Let $I=(x^{4},x^{2}y^{3})$ and $\mathfrak{a}=(x^{2},x^{2}y^{3})$, so that $\mathbf{b}_{1}=(4,0)$, $\mathbf{b}_{2}=(2,3)$, $\mathbf{a}_{1}=(2,0)$, and $\mathbf{a}_{2}=(2,3)$.
Proposition 4.12.
In the current example, the joint Hilbert-Kunz multiplicity as a function of $t$ is given below.
$e_{HK}(\mathfrak{a}^{t};I)=\begin{cases}6t^{2}+12&\text{ if }t<1\\ 3t^{2}+6t+9&\text{ if }t\geq 1\end{cases}$
The Hara-Yoshida $\mathfrak{a}^{t}$-integral closures are given below.
$\overline{I}^{\mathfrak{a}^{t}}=\begin{cases}I_{1}:=\overline{I}=(x^{4},x^{4}y,x^{3}y^{2},x^{2}y^{3})&\text{ if }t<\frac{1}{6}\\ I_{2}:=I_{1}+(x^{3}y)=(x^{4},x^{3}y,x^{3}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{6}\leq t<\frac{1}{3}\\ I_{3}:=I_{2}+(x^{2}y^{2})=(x^{4},x^{3}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{3}\leq t<\frac{1}{2}\\ I_{4}:=I_{3}+(x^{3})=(x^{3},x^{3}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{2}\leq t<\frac{2}{3}\\ I_{5}:=I_{4}+(x^{2}y)=(x^{3},x^{2}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }\frac{2}{3}\leq t<1\\ I_{6}:=I_{5}+(x^{2})=(x^{2},x^{2}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }1\leq t<\frac{7}{6}\\ I_{7}:=I_{6}+(xy)=(x^{2},xy,x^{2}y^{3})&\text{ if }\frac{7}{6}\leq t<\frac{3}{2}\\ I_{8}:=I_{7}+(x)=(x,xy,x^{2}y^{3})&\text{ if }\frac{3}{2}\leq t<2\\ I_{9}:=I_{8}+(1)=(1)&\text{ if }t\geq 2\end{cases}$
The Hara-Yoshida $\mathfrak{a}^{t}$-tight and $\mathfrak{a}^{t}$-Frobenius closures are given below.
$I^{*\mathfrak{a}^{t}}=I^{F\mathfrak{a}^{t}}=\begin{cases}J_{1}:=I=(x^{4},x^{2}y^{3})&\text{ if }t<\frac{1}{2}\\ J_{2}:=J_{1}+(x^{5}y^{2})=(x^{4},x^{5}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{2}\leq t<1\\ J_{3}:=J_{2}+(x^{4}y,x^{4}y^{2})=(x^{4},x^{4}y,x^{4}y^{2},x^{2}y^{3})&\text{ if }1\leq t<\frac{3}{2}\\ J_{4}:=J_{3}+(x^{3},x^{3}y,x^{3}y^{2})=(x^{3},x^{3}y,x^{3}y^{2},x^{2}y^{3})&\text{ if }\frac{3}{2}\leq t<2\\ J_{5}:=J_{4}+(x^{2},x^{2}y,x^{2}y^{2})=(x^{2},x^{2}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }2\leq t<\frac{5}{2}\\ J_{6}:=J_{5}+(x,xy)=(x,xy,x^{2}y^{3})&\text{ if }\frac{5}{2}\leq t<3\\ J_{7}:=J_{6}+(1)=(1)&\text{ if }t\geq 3\end{cases}$
The Vraciu $\mathfrak{a}^{t}$-tight closures are given below.
${}^{\mathfrak{a}^{t}}I^{*}=\begin{cases}K_{1}:=I=(x^{4},x^{2}y^{3})&\text{ if }t<\frac{1}{2}\\ K_{2}:=K_{1}+(x^{5}y^{2})=(x^{4},x^{5}y,x^{2}y^{3})&\text{ if }\frac{1}{2}\leq t<1\\ K_{3}:=K_{2}+(x^{4}y,x^{4}y^{2})=(x^{4},x^{4}y,x^{4}y^{2},x^{2}y^{3})&\text{ if }t\geq 1\end{cases}$
Finally, the Vraciu $\mathfrak{a}^{t}$-Frobenius closures are given below.
${}^{\mathfrak{a}^{t}}I^{F}=\begin{cases}K_{1}=I&\text{ if }t<\frac{1}{2}\\ K_{2}=K_{1}+(x^{5}y^{2})=(x^{4},x^{5}y^{2},x^{2}y^{3})&\text{ if }t\geq\frac{1}{2},\text{ and }t<1\text{ or }p\neq 3\\ K_{3}=K_{2}+(x^{4}y,x^{4}y^{2})=(x^{4},x^{4}y,x^{4}y^{2},x^{2}y^{3})&\text{ if}t\geq 1\text{ and }p=3\end{cases}$
4.3.2. The joint Hilbert-Kunz multiplicity
By Theorem 4.3, we can compute $e_{HK}(\mathfrak{a}^{t};I)$ by finding the area of a certain region in the plane. In the diagrams below, the cone $\sigma^{\vee}$ is bounded by the two rays meeting at the origin; the vertical line segment joining $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$ is $H$, the convex hull of the exponent vectors of the generators of $\mathfrak{a}$, and the other vertical line segments are the sets $\mathbf{b}_{1}+tH$ and $\mathbf{b}_{2}+tH$, for $t<1$ on the left, and for $t\geq 1$ on the right. The shaded regions are the sets $\mathcal{B}+tH+\sigma^{\vee}$. The joint Hilbert-Kunz multiplicity is, therefore, the area of the region inside the cone but outside the shaded region, which can be readily computed from the diagrams.
[…]
4.3.3. Computing the integral closures
A vector $\mathbf{v}$ is in the Newton polytope of $I$ if and only if $\langle\mathbf{c}_{1},\mathbf{v}\rangle\geq 0$, $\langle\mathbf{c}_{2},\mathbf{v}\rangle\geq 0$, and $\langle(3,2),\mathbf{v}\rangle\geq 12$, with the last inequality describing the half-plane with boundary through the vectors $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$. Note that since the generators of $I$ lie on the facets of the cone $\sigma^{\vee}$, a vector in $\sigma^{\vee}$ is on the Newton polytope of $I$ if and only if $\langle(3,2),\mathbf{v}\rangle\geq 12$, equivalently $3v_{1}+2v_{2}\geq 12$.
The integral closure of $I$ is generated by monomials in the Newton polytope, and so we have that, for all $t$,
$\overline{I}={}^{\mathfrak{a}^{t}}\overline{I}=(x^{4},x^{4}y,x^{3}y^{2},x^{2}y^{3}).$
A monomial $x^{v_{1}}y^{v_{2}}$ is in $\overline{I}^{\mathfrak{a}^{t}}$ if and only if $\mathbf{v}+tH$ is contained in the Newton polytope, where $H$ is the convex hull of $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$. Since the Newton polytope is convex, it suffices to check that $\mathbf{v}+t\mathbf{a}_{1}$ and $\mathbf{v}+t\mathbf{a}_{2}$ are in the Newton polytope. Thus, $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$ if and only if the following inequalities hold.
$\langle\mathbf{(}3,2),\mathbf{v}+t\mathbf{a}_{1}\rangle$ $\geq 12$ $\iff$ $3v_{1}+2v_{2}+6t$ $\geq 12$
$\langle\mathbf{(}3,2),\mathbf{v}+t\mathbf{a}_{2}\rangle$ $\geq 12$ $\iff$ $3v_{1}+2v_{2}+12t$ $\geq 12$
Therefore $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$ if and only if $3v_{1}+2v_{2}+6t\geq 12$, i.e. $t\geq\frac{12-3v_{1}-2v_{2}}{6}$. For each of the eight monomials in $R$ outside $\overline{I}$, we can compute the value of $t$ which puts that monomial into $\overline{I}^{\mathfrak{a}^{t}}$ and describe this ideal for all $t$.
$\begin{array}[]{|c|c|c|}\hline\cr x^{\mathbf{v}}&12-3v_{1}-2v_{2}&t\text{ suchthat }x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}\\ \hline\cr x^{3}y&1&t\geq 1/6\\ x^{2}y^{2}&2&t\geq 1/3\\ x^{3}&3&t\geq 1/2\\ x^{2}y&4&t\geq 2/3\\ x^{2}&6&t\geq 1\\ xy&7&t\geq 7/6\\ x&9&t\geq 3/2\\ 1&12&t\geq 2\\ \hline\cr\end{array}$
Therefore,
$\overline{I}^{\mathfrak{a}^{t}}=\begin{cases}I_{1}:=\overline{I}=(x^{4},x^{4}y,x^{3}y^{2},x^{2}y^{3})&\text{ if }t<\frac{1}{6}\\ I_{2}:=I_{1}+(x^{3}y)=(x^{4},x^{3}y,x^{3}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{6}\leq t<\frac{1}{3}\\ I_{3}:=I_{2}+(x^{2}y^{2})=(x^{4},x^{3}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{3}\leq t<\frac{1}{2}\\ I_{4}:=I_{3}+(x^{3})=(x^{3},x^{3}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{2}\leq t<\frac{2}{3}\\ I_{5}:=I_{4}+(x^{2}y)=(x^{3},x^{2}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }\frac{2}{3}\leq t<1\\ I_{6}:=I_{5}+(x^{2})=(x^{2},x^{2}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }1\leq t<\frac{7}{6}\\ I_{7}:=I_{6}+(xy)=(x^{2},xy,x^{2}y^{3})&\text{ if }\frac{7}{6}\leq t<\frac{3}{2}\\ I_{8}:=I_{7}+(x)=(x,xy,x^{2}y^{3})&\text{ if }\frac{3}{2}\leq t<2\\ I_{9}:=I_{8}+(1)=(1)&\text{ if }t\geq 2\end{cases}$
4.3.4. Computing the tight and Frobenius closures
We will use Theorem 4.10. We have that $\mathbf{a}_{2}-\mathbf{a}_{1}=(0,3)$, and so $\langle\mathbf{c}_{1},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle=3$ and $\langle\mathbf{c}_{2},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle=-6$. Therefore, for any $\mathbf{v}=(v_{1},v_{2})\in S$, we have that
$m_{i}(\mathbf{v})=\frac{\langle\mathbf{c}_{1},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{1},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{1},\mathbf{a}_{1}\rangle}{\langle\mathbf{c}_{1},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle}=\frac{\langle\mathbf{c}_{1},\mathbf{b}_{i}\rangle-v_{2}}{3}=\begin{cases}-\frac{v_{2}}{3}&\text{if }i=1\\ \frac{3-v_{2}}{3}&\text{if }i=2\end{cases}=\frac{3i-3-v_{2}}{3}$
and
$M_{i}(\mathbf{v})=\frac{\langle\mathbf{c}_{2},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{2},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{2},\mathbf{a}_{1}\rangle}{\langle\mathbf{c}_{2},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle}=\frac{\langle\mathbf{c}_{2},\mathbf{b}_{i}\rangle-(3v_{1}-2v_{2})-6(t-s)}{-6}$
=$=\begin{cases}-\frac{1}{6}(12-3v_{1}+2v_{2}-6(t-s))&\text{if }i=1\\ -\frac{1}{6}(-3v_{1}+2v_{2}-6(t-s))&\text{if }i=2\end{cases}=\frac{12i-24+3v_{1}-2v_{2}}{6}+t-s$
For $i=1,2$, let
$U_{i}=\begin{cases}\left[\frac{m_{i}(\mathbf{v})}{t},\frac{M_{i}(\mathbf{v})+s}{t}\right]&\text{ if }m_{i}(\mathbf{v})\leq M_{i}(\mathbf{v})\\ \varnothing&\text{ otherwise}\end{cases}$
By Theorem 4.10 we have that $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ if and only if $[0,1]\subset U_{1}\cup U_{2}$. Now,
$m_{i}(\mathbf{v})\leq M_{i}(\mathbf{v})\iff\frac{3i-3-v_{2}}{3}\leq\frac{12i-24+3v_{1}-2v_{2}}{6}+t-s\iff v_{1}\geq 6-2i-2(t-s)$
in which case
$U_{i}=\left[\frac{3i-3-v_{2}}{3t},\frac{12i-24+3v_{1}-2v_{2}}{6t}+1\right].$
Suppose that $[0,1]\subset U_{1}$. This implies that
$\frac{-v_{2}}{3t}$ $\leq 0$$\implies$ $v_{2}$ $\geq 0$
$\frac{-12+3v_{1}-2v_{2}}{6t}+1$ $\geq 1$$\implies$ $3v_{1}-2v_{2}$ $\geq 12.$
In particular, this means that $\langle\mathbf{c}_{1},\mathbf{v}-\mathbf{b}_{1}\rangle=v_{2}\geq 0$ and $\langle\mathbf{c}_{2},\mathbf{v}-\mathbf{b}_{1}\rangle=3v_{1}-2v_{2}-12\geq 0$. Therefore $\mathbf{v}-\mathbf{b}_{1}\in\sigma^{\vee}$, implying that $\mathbf{v}\in\mathbf{b}_{1}+\sigma^{\vee}$ and so $x^{\mathbf{v}}\in I$.
Similarly, suppose that $[0,1]\subset U_{2}$, so that
$\frac{3-v_{2}}{3t}$ $\leq 0$$\implies$ $v_{2}$ $\geq 3$
$\frac{3v_{1}-2v_{2}}{6t}+1$ $\geq 1$$\implies$ $3v_{1}-2v_{2}$ $\geq 0.$
Therefore $\langle\mathbf{c}_{1},\mathbf{v}-\mathbf{b}_{2}\rangle=v_{2}-3\geq 0$ and $\langle\mathbf{c}_{2},\mathbf{v}-\mathbf{b}_{2}\rangle=3v_{1}-2v_{2}\geq 0$. Therefore $\mathbf{v}-\mathbf{b}_{2}\in\sigma^{\vee}$, implying that $\mathbf{v}\in\mathbf{b}_{2}+\sigma^{\vee}$ and so $x^{\mathbf{v}}\in I$.
Therefore, the only situation in which $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ but $x^{\mathbf{v}}\notin I$ is when $[0,1]$ is contained in the union of $U_{1}$ and $U_{2}$ but is contained in neither $U_{1}$ and $U_{2}$. For this situation to occur we must have that
[…]
This system simplifies to
$v_{1}$ $\geq 4-2(t-s)$
$v_{2}$ $\geq 0$
$v_{1}$ $\geq 6-2t$
$3v_{1}-2v_{2}$ $\geq 0$
The second and fourth inequalities above are satisfied for all $\mathbf{v}\in S$. Thus, we have that for $\mathbf{v}\in S$, $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ if and only if $x^{\mathbf{v}}\in I$ or $v_{1}\geq\max\{4-2(t-s),6-2t\}$.
4.3.5. The Hara-Yoshida tight and Frobenius closures
By Corollary 4.6, $x^{\mathbf{v}}\in I^{*\mathfrak{a}^{t}}=I^{F\mathfrak{a}^{t}}$ if and only if $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$. By the analysis above, and taking $s=0$, this is satisfied for any $\mathbf{v}$ with $x^{\mathbf{v}}\in I$ or $v_{1}\geq\max\{4-2(t-0),6-2t\}=6-2t$. So, we have that $I^{*\mathfrak{a}^{t}}=I+(x^{\mathbf{v}}\mid\mathbf{v}\in S,v_{1}\geq 6-2t)$. Therefore,
$I^{*\mathfrak{a}^{t}}=I^{F\mathfrak{a}^{t}}=\begin{cases}J_{1}:=I=(x^{4},x^{2}y^{3})&\text{ if }t<\frac{1}{2}\\ J_{2}:=J_{1}+(x^{5}y^{2})=(x^{4},x^{5}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{2}\leq t<1\\ J_{3}:=J_{2}+(x^{4}y,x^{4}y^{2})=(x^{4},x^{4}y,x^{4}y^{2},x^{2}y^{3})&\text{ if }1\leq t<\frac{3}{2}\\ J_{4}:=J_{3}+(x^{3},x^{3}y,x^{3}y^{2})=(x^{3},x^{3}y,x^{3}y^{2},x^{2}y^{3})&\text{ if }\frac{3}{2}\leq t<2\\ J_{5}:=J_{4}+(x^{2},x^{2}y,x^{2}y^{2})=(x^{2},x^{2}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }2\leq t<\frac{5}{2}\\ J_{6}:=J_{5}+(x,xy)=(x,xy,x^{2}y^{3})&\text{ if }\frac{5}{2}\leq t<3\\ J_{7}:=J_{6}+(1)=(1)&\text{ if }t\geq 3\end{cases}$
4.3.6. The Vraciu tight closure
By Theorem 4.5, $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}I^{*}$ if and only if $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$. By the analysis above, and taking $s=t$, this is satisfied for any $\mathbf{v}$ with $x^{\mathbf{v}}\in I$ or $v_{1}\geq\max\{4-2(t-t),6-2t\}=\max\{4,6-2t\}$. Therefore,
${}^{\mathfrak{a}^{t}}I^{*}=\begin{cases}K_{1}:=I=(x^{4},x^{2}y^{3})&\text{ if }t<\frac{1}{2}\\ K_{2}:=K_{1}+(x^{5}y^{2})=(x^{4},x^{5}y,x^{2}y^{3})&\text{ if }\frac{1}{2}\leq t<1\\ K_{3}:=K_{2}+(x^{4}y,x^{4}y^{2})=(x^{4},x^{4}y,x^{4}y^{2},x^{2}y^{3})&\text{ if }t\geq 1\end{cases}$
4.3.7. The Vraciu Frobenius closure
By Theorem 4.5, a sufficient condition for $x^{\mathbf{v}}$ to be in ${}^{\mathfrak{a}^{t}}I^{F}$ is for $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ for some $s>t$. Letting $s=t+\epsilon$ for some small $\epsilon>0$, we have that this condition is satisfied for any $\mathbf{v}$ with $x^{\mathbf{v}}\in I$ or $v_{1}\geq\max\{4+2\epsilon,6-2t\}$. Thus, we have that if $t\geq\frac{1}{2}$, then ${}^{\mathfrak{a}^{t}}I^{F}\supseteq I+(x^{5}y^{2})$.
Now, if $t<\frac{1}{2}$, then $I\subset{}^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{t}}I^{*}=I$, so ${}^{\mathfrak{a}^{t}}I^{F}=I$. If $\frac{1}{2}\leq t<1$, then $I+(x^{5}y^{2})\subset{}^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{t}}I^{*}=I+(x^{5}y^{2})$, and so ${}^{\mathfrak{a}^{t}}I^{F}=I+(x^{5}y^{2})$. If $t\geq 1$, then $I+(x^{5}y^{2})\subset{}^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{t}}I^{*}=I+(x^{4}y,x^{4}y^{2})$.
To determine ${}^{\mathfrak{a}^{t}}I^{F}$ for $t\geq 1$, we use the definition to check whether the monomials $x^{4}y$ and $x^{4}y^{2}$ belong to ${}^{\mathfrak{a}^{t}}I^{F}$. For a power $q$ of $p$ we have that
$(x^{4}y)^{q}\mathfrak{a}^{\lceil tq\rceil}=(x^{4}y)^{q}(x^{2},x^{2}y^{3})^{\lceil tq\rceil}=(x^{4q+2\lceil tq\rceil}y^{q+3i}\mid 0\leq i\leq\lceil tq\rceil)$
and
$\mathfrak{a}^{\lceil tq\rceil}I^{[q]}=(x^{2},x^{2}y^{3})^{\lceil tq\rceil}(x^{4q},x^{2q}y^{3q})=(x^{4q+2\lceil tq\rceil}y^{3j},x^{2q+2\lceil tq\rceil}y^{3q+3j}\mid 0\leq j\leq\lceil tq\rceil).$
Now $(x^{4}y)^{q}\mathfrak{a}^{\lceil tq\rceil}\subset\mathfrak{a}^{\lceil tq\rceil}I^{[q]}$ if and only if for every $i$, $0\leq i\leq\lceil tq\rceil$, there exists $j$, $0\leq j\leq\lceil tq\rceil$, such that $x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{4q+\lceil tq\rceil}y^{3j})$ or $x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{2q+\lceil tq\rceil}y^{3q+3j})$. Now, in $R$, we have that $x^{a}y^{b}\in(x^{c}y^{d})$ if and only if $(a,b)-(c,d)\in S$, which occurs exactly when $0\leq 2(b-d)\leq 3(a-c)$. Therefore,
$x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{4q+2\lceil tq\rceil}y^{3j})\iff 0\leq 2(q+3i-3j)\leq 3(4q+2\lceil tq\rceil-(4q+2\lceil tq\rceil))=0.$
Thus, $x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{4q+2\lceil tq\rceil}y^{3j})$ if and only if $q+3i-3j=0$. This is impossible unless $p=3$, in which case $j=\frac{q+3i}{3}$. Since $j\leq\lceil tq\rceil$, this is satisfiable so long as $q+3i\leq 3\lceil tq\rceil$, that is, $i\leq\lceil tq\rceil-\frac{q}{3}$.
On the other hand,
$x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{2q+2\lceil tq\rceil}y^{3q+3j})\iff 0\leq 2(q+3i-(3q+3j))\leq 3(4q+2\lceil tq\rceil-(2q+2\lceil tq\rceil)),$
which simplifies to
$0\leq 6i-6j-4q\leq 6q\implies i-\frac{5}{3}q\leq j\leq i-\frac{2}{3}q.$
There exists a $j$ satisfying this so long as $i-\frac{2}{3}q\geq 0$, since $i-\frac{5}{3}q\leq i\leq\lceil tq\rceil$.
Thus, we have that $x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{4q+2\lceil tq\rceil}y^{3j})$ for some $j$ if and only if $p=3$ and $i\leq\lceil tq\rceil-\frac{q}{3}$, and we have that $x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{2q+2\lceil tq\rceil}y^{3q+3j})$ for some $j$ if and only if $i\geq\frac{2}{3}q$. In order for at least one of these containments to hold for each $0\leq i\leq\lceil tq\rceil$, we must have that $p=3$ and $\frac{2}{3}q\leq\lceil tq\rceil-\frac{q}{3}$, i.e. $q\leq\lceil tq\rceil$. This holds for all $q\gg 0$ if and only if $t\geq 1$.
Working similarly to check $x^{4}y^{2}$, we have that
$(x^{4}y^{2})^{q}\mathfrak{a}^{\lceil tq\rceil}=(x^{4}y^{2})^{q}(x^{2},x^{2}y^{3})^{\lceil tq\rceil}=(x^{4q+2\lceil tq\rceil}y^{2q+3i}\mid 0\leq i\leq\lceil tq\rceil)$
and as before
$\mathfrak{a}^{\lceil tq\rceil}I^{[q]}=(x^{4q+2\lceil tq\rceil}y^{3j},x^{2q+2\lceil tq\rceil}y^{3q+3j}\mid 0\leq j\leq\lceil tq\rceil).$
Now $(x^{4}y^{2})^{q}\mathfrak{a}^{\lceil tq\rceil}\subset\mathfrak{a}^{\lceil tq\rceil}I^{[q]}$ if and only if for every $i$, $0\leq i\leq\lceil tq\rceil$, there exists $j$, $0\leq j\leq\lceil tq\rceil$, such that $x^{4q+2\lceil tq\rceil}y^{2q+3i}\in(x^{4q+\lceil tq\rceil}y^{3j})$ or $x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{2q+\lceil tq\rceil}y^{3q+3j})$. We have that
$x^{4q+2\lceil tq\rceil}y^{2q+3i}\in(x^{4q+2\lceil tq\rceil}y^{3j})\iff 0\leq 2(2q+3i-3j)\leq 3(4q+2\lceil tq\rceil-(4q+2\lceil tq\rceil))=0.$
Thus, $x^{4q+2\lceil tq\rceil}y^{2q+2i}\in(x^{4q+2\lceil tq\rceil}y^{2j})$ if and only if $02q+3i-3j=$ 0, which implies $j=\frac{2q+3i}{3}$. This is again only possible if $p=3$, in which case, we can find such a $j$ so long as $2q+3i\leq 3\lceil tq\rceil$, that is, $i\leq\lceil tq\rceil-\frac{2}{3}q$.
On the other hand,
$x^{4q+2\lceil tq\rceil}y^{2q+3i}\in(x^{2q+2\lceil tq\rceil}y^{3q+3j})\iff 0\leq 2(2q+3i-(3q+3j))\leq 3(4q+2\lceil tq\rceil-(2q+2\lceil tq\rceil)),$
which simplifies to
$0\leq 6i-6j-2q\leq 6q\implies i-\frac{4}{3}q\leq j\leq i-\frac{1}{3}q.$
There exists a $j$ satisfying this so long as $i-\frac{1}{3}q\geq 0$, since $i-\frac{4}{3}q\leq i\leq\lceil tq\rceil$.
Thus, we have that $x^{4q+2\lceil tq\rceil}y^{2q+3i}\in(x^{4q+2\lceil tq\rceil}y^{3j})$ for some $j$ if and only if $p=3$ and $i\leq\lceil tq\rceil-\frac{2}{3}q$, and we have that $x^{4q+2\lceil tq\rceil}y^{2q+3i}\in(x^{2q+2\lceil tq\rceil}y^{3q+3j})$ for some $j$ if and only if $i\geq\frac{1}{3}q$. In order for at least one of these containments to hold for each $0\leq i\leq\lceil tq\rceil$, we must have that $p=3$ and $\frac{1}{3}q\leq\lceil tq\rceil-\frac{2}{3}q$, i.e. $q\leq\lceil tq\rceil$. This holds for all $q\gg 0$ if and only if $t\geq 1$.
So, we have found that ${}^{\mathfrak{a}^{t}}I^{F}$ for $t\geq 1$ depends on $p$, and hence cannot be detected simply using the convex geometry of the affine semigroup. Finally,
${}^{\mathfrak{a}^{t}}I^{F}=\begin{cases}K_{1}=I&\text{ if }t<\frac{1}{2}\\ K_{2}=K_{1}+(x^{5}y^{2})=(x^{4},x^{5}y^{2},x^{2}y^{3})&\text{ if }t\geq\frac{1}{2},\text{ and }t<1\text{ or }p\neq 3\\ K_{3}=K_{2}+(x^{4}y,x^{4}y^{2})=(x^{4},x^{4}y,x^{4}y^{2},x^{2}y^{3})&\text{ if}t\geq 1\text{ and }p=3\end{cases}$
4.4. A 3-generated example
Let $R=k[x,y]$, $\mathfrak{a}=(xy^{3},x^{2}y,x^{3})$, and $I=(x^{4}y,xy^{4})$, so that $\mathbf{b}_{i}=(7-3i,3i-2)$ for $i=1,2$. We wish to compute the various $\mathfrak{a}^{t}$-closures of $I$. In this case we have that $\Delta=\{(\alpha,\beta,\gamma)^{T}\mid\alpha+\beta+\gamma=1,\alpha,\beta,\gamma\geq 0\}$.
For $\mathbf{u}=(\alpha,\beta,\gamma)^{T}\in\Delta$ and $\mathbf{v}=(v_{1},v_{2})^{T}$, we have that
$\mathbf{v}+tA\mathbf{u}\in\mathbf{\mathcal{missing}}{B}+\sigma^{\vee}$
$\iff\exists i\in\{1,2\}:\quad C^{T}(\mathbf{v}-\mathbf{b}_{i}+tA\mathbf{u})\geq\mathbf{0}$
$\iff\exists i\in\{1,2\}:\quad\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)\left(\left(\begin{array}[]{c}v_{1}\\ v_{2}\end{array}\right)-\left(\begin{array}[]{c}7-3i\\ 3i-2\end{array}\right)+t\left(\begin{array}[]{ccc}1&2&3\\ 3&1&0\end{array}\right)\left(\begin{array}[]{c}\alpha\\ \beta\\ \gamma\end{array}\right)\right)\geq\mathbf{0}$
$\iff\exists i\in\{1,2\}:\quad\left(\begin{array}[]{c}v_{2}-3i+2\\ v_{1}-7+3i\end{array}\right)+t\left(\begin{array}[]{c}3\alpha+\beta\\ \alpha+2\beta+3\gamma\end{array}\right)\geq\mathbf{0}$
$\iff\exists i\in\{1,2\}:\quad t\left(\begin{array}[]{c}3\alpha+\beta\\ \alpha+2\beta+3\gamma\end{array}\right)\geq\left(\begin{array}[]{c}3i-2-v_{2}\\ 7-3i-v_{1}\end{array}\right)$
$\iff\exists i\in\{1,2\}:\quad t\geq\max\left(\frac{3i-2-v_{2}}{3\alpha+\beta},\frac{7-3i-v_{1}}{\alpha+2\beta+3\gamma}\right)$
$\iff t\geq\min\left(\max\left(\frac{1-v_{2}}{3\alpha+\beta},\frac{4-v_{1}}{\alpha+2\beta+3\gamma}\right),\max\left(\frac{4-v_{2}}{3\alpha+\beta},\frac{1-v_{1}}{\alpha+2\beta+3\gamma}\right)\right),$
where if $03\alpha+\beta=$ 0, we interpret $\frac{3i-2-v_{2}}{3\alpha+\beta}$ as $\infty$ if $03i-2-v_{2}>$ 0, $\infty$- ∞ if $03i-2-v_{2}<$ 0, and 0 if $03i-2-v_{2}=$ 0. Thus, $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$ if and only if for all $(\alpha,\beta,\gamma)^{T}\in\Delta$, the above inequality holds, that is,
$t\geq\sup_{(\alpha,\beta\gamma)^{T}\in\Delta}\left\{\min\left(\max\left(\frac{1-v_{2}}{3\alpha+\beta},\frac{4-v_{1}}{\alpha+2\beta+3\gamma}\right),\max\left(\frac{4-v_{2}}{3\alpha+\beta},\frac{1-v_{1}}{\alpha+2\beta+3\gamma}\right)\right)\right\}.$
By way of example, we find the smallest $t$ such that $x^{3}y^{2}\in I^{F\mathfrak{a}^{t}}$. This gives us $v_{1}=3$, $v_{2}=2$, and so we wish to compute
$\sup_{(\alpha,\beta\gamma)^{T}\in\Delta}\left\{\min\left(\max\left(\frac{1-v_{2}}{3\alpha+\beta},\frac{4-v_{1}}{\alpha+2\beta+3\gamma}\right),\max\left(\frac{4-v_{2}}{3\alpha+\beta},\frac{1-v_{1}}{\alpha+2\beta+3\gamma}\right)\right)\right\}$
$=$ $\sup_{(\alpha,\beta\gamma)^{T}\in\Delta}\left\{\min\left(\max\left(\frac{-1}{3\alpha+\beta},\frac{1}{\alpha+2\beta+3\gamma}\right),\max\left(\frac{2}{3\alpha+\beta},\frac{-2}{\alpha+2\beta+3\gamma}\right)\right)\right\}$
$=$ $\sup_{(\alpha,\beta\gamma)^{T}\in\Delta}\left\{\min\left(\frac{1}{\alpha+2\beta+3\gamma},\frac{2}{3\alpha+\beta}\right)\right\}.$
Now $\frac{2}{3\alpha+\beta}\leq\frac{1}{\alpha+2\beta+3\gamma}$ if and only if $\alpha-3\beta-6\gamma\geq 0$, and so the supremum above is equal to
$\sup\left\{\sup_{\alpha-3\beta-6\gamma\geq 0}\left\{\frac{2}{3\alpha+\beta}\right\},\sup_{\alpha-3\beta-6\gamma\leq 0}\left\{\frac{1}{\alpha+2\beta+3\gamma}\right\}\right\}=\sup\left\{\frac{4}{5},\frac{4}{5}\right\}=\frac{4}{5}.$
Therefore $x^{3}y^{2}\in I^{F\mathfrak{a}^{t}}$ so long as $t\geq\frac{4}{5}$.